### v1

# Doubly-Robust LLM-as-a-Judge: Externally Valid Estimation with Imperfect Personas

Abstract
As Generative AI (GenAI) systems see growing adoption, a key concern involves the external validity of evaluations, or the extent to which they generalize from lab-based to real-world deployment conditions. Threats to the external validity of GenAI evaluations arise when the source sample of human raters and system outputs used to obtain a system quality estimate differs from the target distribution at deployment time. In this work, we propose a doubly-robust estimation framework designed to address this evaluation sampling bias. Key to our approach is the use of “persona” ratings produced by prompting an LLM evaluator (i.e., an LLM-as-a-judge) to behave as a human rater with specific sociodemographic characteristics. Our doubly-robust framework combines these informative yet imperfect persona ratings with human ratings obtained under evaluation sampling bias to produce statistically valid system quality estimates. In particular, we show that our approach yields valid system quality estimates when either (i) a model trained to predict human ratings using persona ratings and source data observed under sampling bias, or (ii) a reweighting model that corrects for sampling bias is of sufficient quality. We validate our framework theoretically and via a novel Persona Simulation Framework (PSF) designed to systematically manipulate persona quality and the degree of evaluation sampling bias present in source data. Our work provides a principled foundation for combining imperfect persona ratings with human ratings observed under sampling bias to obtain valid system quality estimates.
1 Introduction
As Generative AI (GenAI) systems see growing adoption, a key concern involves the external validity of evaluations, or the extent to which they generalize from lab-based to real-world deployment conditions (Ibrahim et al., 2024; Ouyang et al., 2023; Liao & Xiao, 2023; Liao et al., 2021; Weidinger et al., 2025). In particular, many evaluations report a system quality estimate, which reflects the proportion of outputs rated to exhibit a capability (e.g., “helpfulness”) or defect (e.g., “toxicity”) by a human with specialized characteristics (e.g., domain knowledge, cultural experience). However, such quality estimates may fail to generalize when the source distribution of human raters and system outputs available at evaluation time differs from the target distribution encountered upon deployment. For example, in medical visit summarization, covariate shift arises when we wish to obtain a system quality estimate via expert physician ratings, but rely primarily on medical student ratings at evaluation time (Cai et al., 2022). Selection bias may also occur if medical students choose to rate complex summaries less frequently than more experienced physicians. Left unaddressed, these forms of evaluation sampling bias threaten the external validity of system quality estimates.
Recent work has proposed tools for improving system quality estimates when human ratings are scarce but black-box predictions (e.g., from an LLM-as-a-judge) are cheap and abundant (Chatzi et al., 2024; Fisch et al., 2024; Eyre & Madras, 2024; Dorner et al., 2024; Saad-Falcon et al., 2023; Fogliato et al., 2024). For instance, Prediction Powered Inference (PPI) offers an approach for leveraging a subset of labeled (source) data to correct for bias in black-box model predictions (Angelopoulos et al., 2023a; b). This bias correction enables using black-box predictions generated over unlabeled (target) samples to shrink confidence intervals around quality estimates while maintaining valid coverage. However, both PPI and its extensions (e.g., PPI++ (Angelopoulos et al., 2023b), RePPI (Ji et al., 2025)) assume that (i) source and target samples are drawn i.i.d. and (ii) labels are missing completely at random (MCAR) (Tsiatis, 2006), i.e. that successful completion of a rating is independent of rater and text characteristics. However, when source data is observed under sampling bias, these assumptions are violated and severe miscoverage occurs (see Fig. 1).
In this work, we devise an estimator that directly corrects for evaluation sampling bias. Like existing approaches (Angelopoulos et al., 2023b; Ji et al., 2025), our proposal leverages black-box predictions generated by a GenAI system over unlabeled (target) samples to improve statistical inference. Unlike existing estimators, however, our proposal is doubly-robust (Bang & Robins, 2005; Chernozhukov et al., 2018): it yields valid system quality estimates if either a model trained to predict human ratings from source samples or a reweighting model that corrects for sampling bias is of sufficient quality. To attain this doubly-robust property, we leverage persona ratings — scores generated by prompting an LLM-as-a-judge to behave as a human rater with desired characteristics (e.g., demographics, expertise) — to learn a high-quality predictor for human ratings from labeled source data. This novel perspective treats persona ratings as an informative yet imperfect proxy for human ratings to enhance the quality of downstream system quality estimates.
Figure 1 illustrates the benefits of our estimation approach on three datasets. Whereas directly adopting persona ratings for estimation (Persona-Based) and state-of-the-art baselines (Angelopoulos et al., 2023b; Ji et al., 2025) fail to provide coverage, our approach provides valid confidence intervals with low bias. To summarize, our main contributions are:
Figure 1: Comparison of our doubly-robust estimator with baselines on three datasets from our Persona Simulation Framework. Black and red dashed lines denote the true source and target mean ratings, respectively (e.g., the average “helpfulness” rating obtained over source vs. target distributions). Baselines yield invalid coverage, while ours achieves valid coverage and low bias.
  • We formalize the problem of GenAI system quality estimation under evaluation sampling bias. Unlike existing statistical frameworks (Angelopoulos et al., 2023b; Ji et al., 2025), our formulation explicitly accounts for both covariate shift and selection bias in observed source ratings to improve the external validity of system quality estimates.
  • We devise a doubly-robust estimator for GenAI system quality estimation under evaluation sampling bias. En route, we first advance doubly-robust estimation theory by generalizing the work of (Chernozhukov et al., 2023) to M-estimation settings with surrogate (persona) ratings. This generalization enables us to (i) leverage persona ratings to improve doubly-robust system quality estimates, (ii) estimate a richer set of system quality parameters (e.g., rating variance, quantiles) beyond means, and (iii) maintain valid coverage even in the presence of evaluation sampling bias, all desiderata not satisfied by previous works.
  • We advance the practical application of doubly-robust estimators to GenAI system quality estimation. Whereas doubly-robust estimators are traditionally applied on small tabular datasets, GenAI system quality estimation requires learning a reweighting function over high dimensional (e.g., text, audio) input-output spaces. We show that sentence transformer embedding models and a “Riesz loss” approach (Chernozhukov et al., 2022b) can be combined to correct for covariate shift in high-dimensional text input-output spaces.
  • We introduce a Persona Simulation Framework (PSF) that systematically manipulates evaluation sampling bias and persona quality over three datasets encompassing synthetic, semi-synthetic (PRISM) (Kirk et al., 2024), and real-world (DICES) (Aroyo et al., 2023) settings. Leveraging the PSF, we show our estimator obtains valid coverage and low bias up to a larger magnitude of sampling bias than state-of-the-art baselines (e.g., RePPI (Ji et al., 2025)).
[…]
2 Preliminaries
We provide an overview of the data generating processes and GenAI system quality parameters considered hereinafter. We provide detailed coverage of all notation and assumptions in Appendix B.
[…]
Example 1.
Suppose we want to measure the “factual consistency” of medical visit summarization model outputs. Here, $X$ captures raters’ expertise (medical student vs. physician) and $V$ includes the full visit notes and corresponding summary. Rating completion ($C$) denotes whether the rater successfully provides a rating when prompted, and $Y$ is the human’s “factual consistency” rating (e.g., on a binary scale). Finally, $\hat{Y}$ denotes GPT-4’s predicted rating.
[…]
While we relax this MCAR assumption, we rely on several additional assumptions (also required by existing frameworks). For instance, we assume no concept drift, i.e., that $P_{s}(Y|W)=P_{t}(Y|W)$. This requires that the rater and content characteristics are sufficiently rich as to describe ratings across both populations. We elaborate on this and other standard causal assumptions required by our framework in Appendix B. Relaxing these assumptions remains a fruitful direction for future work.
Estimation Goal. Given a sample of $N_{s}$ partial source observations $\mathcal{D}_{s}=\{(X^{s}_{j},V^{s}_{j},C^{s}_{j},C^{s}_{j}\cdot Y^{s}_{j},\hat{Y}^{s}_{j})\}_{j=1}^{N_{s}}$ and $N_{t}$ partial target observations $\mathcal{D}_{t}=\{(X^{t}_{i},V^{t}_{i},\hat{Y}^{t}_{i})\}_{i=1}^{N_{t}}$, our goal is to estimate a parameter summarizing system quality over $P_{t}$. [^1] We use superscripts $s$, $t$ on random variables to denote source and target membership. We omit these superscripts where the distribution is clear from context (e.g., $\mathbb{E}_{t}[Y]$ clearly refers to the target distribution). [^2] In the tuple $\mathcal{D}_{s}$, the shorthand $C^{s}_{j}\cdot Y^{s}_{j}$ denotes that ratings are only observed when $C^{s}_{j}=1$. As our running example in the main text and experiments, we consider the mean rating, $\theta_{t}:=\mathbb{E}_{t}[Y]$, which describes the average “helpfulness” or “factual consistency” rating assigned by human raters to system outputs in the target distribution $P_{t}$.
3 Methodology
We now introduce our doubly-robust estimator for GenAI system quality estimation under evaluation sampling bias. The central challenge addressed by our approach is that our data is imperfect. While persona predictions are available, they may be a poor proxy for human ratings. Likewise, human ratings from the source distribution may suffer from evaluation sampling bias, leading to invalid estimates for the target parameter. We first introduce several naive approaches which might be used to tackle this problem ($\S$ 3.1). Then, we show that while each approach is insufficient in isolation, they can be combined to obtain valid coverage. ($\S$ 3.2). Our results presented in this section apply not only when $\theta_{t}:=\mathbb{E}_{t}[Y]$ (where they generalize Chernozhukov et al. (2023) to simultaneous covariate shift and selection bias) but also when $\theta_{t}$ is the solution to a generic M-estimation problem (e.q. (5)).
3.1 Baseline Approaches and Their Limitations
[…]
Re-weighting. Another approach is to disregard the persona ratings. One instead might re-weight samples from $P_{s}$ based on their probability of occurring under $P_{t}$. This approach requires correcting for covariate shift and selection bias in parallel. Formally, let $\omega_{0}(w)=\frac{dP_{t}}{dP_{s}}(w)$ denote the density ratio between $P_{t}(W)$ and $P_{s}(W)$, and let $\pi_{0}(w)=\mathbb{P}_{s}(C=1\mid W=w)$ denote the probability of rating completion. Under standard assumptions (see Appendix B), we have $\theta_{t}=\mathbb{E}_{s}[\alpha_{0}(W,C)Y]$, where $\alpha_{0}(W,C):=C\frac{\omega_{0}(W)}{\pi_{0}(W)}$. Thus, if one produces an ML estimate $\widehat{\alpha}$ of $\alpha_{0}$ (say by training models $\widehat{\omega}$, $\widehat{\pi}$ predicting $\omega_{0}$, $\pi_{0}$), they can compute an inverse propensity weighted (IPW) estimate $\hat{\theta}^{\text{ipw}}_{t}:=\frac{1}{N_{s}}\sum_{j=1}^{N_{s}}\hat{\alpha}(W^{s}_{j},C^{s}_{j})\cdot Y^{s}_{j}$. Again, estimates of $\alpha_{0}$ must converge at parametric rates in order to maintain coverage. Further, IPW suffers from high variance when propensities are small — a salient challenge when estimating system quality parameters over high-dimensional (e.g., text) data ($\S$ 4).
3.2 Doubly-Robust Estimator
Our doubly-robust estimator can be viewed as carefully combining the persona-augmented regression estimator ($\mu_{0}$) with the re-weighting estimator ($\alpha_{0}$). The functions $\mu_{0}$ and $\alpha_{0}$ are referred to as nuisance functions because they are used as an auxiliary information source to estimate the target statistical parameter of interest $\theta_{t}$. Our estimator combines these nuisance functions in the form:
$\widehat{\theta}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\mu}(W_{i}^{t},\widehat{Y}_{i}^{t})+\frac{1}{N_{s}}\sum_{j=1}^{N_{s}}\widehat{\alpha}(W_{j}^{s},C_{j}^{s})\left\{Y_{j}^{s}-\hat{\mu}(W_{j}^{s},\widehat{Y}_{j}^{s})\right\},$ (1)
where $\widehat{\mu}$ and $\widehat{\alpha}$ are estimates of $\mu_{0}$ and $\alpha_{0}$ that are assumed to be independent of the data.
In Equation (1), the left term evaluates the regression-based estimator over samples from $P_{t}$. The right term corrects for bias in the human rating predictor $\hat{\mu}$ by re-weighting residualized source data to account for covariate shift and selection bias. To construct confidence intervals, we also consider the variance estimate:
$\hat{\sigma}^{2}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\left\{\widehat{\mu}(W_{i}^{t},\hat{Y}_{i}^{t})-\widehat{m}_{t}\right\}^{2}+\frac{\widehat{\gamma}}{N_{s}}\sum_{j=1}^{N_{s}}\widehat{\alpha}(W_{j}^{s},C_{j}^{s})^{2}\left\{Y_{j}^{s}-\widehat{\mu}(W_{j}^{s},\hat{Y}_{j}^{s})\right\}^{2},$ (2)
where $\widehat{m}_{t}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\mu}(W_{i}^{t},\hat{Y}_{i}^{t})$ and $\widehat{\gamma}$ is a scaling parameter (described in Algorithm 1). Since our mean and variance estimators require nuisance estimates that are independent of the data, we use $K$-fold cross-fitting to maximize efficiency (Chernozhukov et al., 2018). For each $k\leq K$, we train nuisance models on all data excluding the samples in fold $k$. We then use our nuisance estimates to produce a de-biased parameter estimate for the data in fold $k$. We then average the per-fold parameter and variance estimates to maintain full data efficiency. See Algorithms 1 and 2 for full details.
Our main result establishes the asymptotic normality of our estimator. Further, it describes how to build confidence intervals using the mean and variance estimates recovered from Algorithm 1.
Algorithm 1 Doubly-Robust Estimator with $K$-fold Cross-Fitting
1:Input: Samples $\mathcal{D}_{s}=\{Z_{1}^{s},\dots,Z_{N_{s}}^{s}\}$ from $P_{s}$, samples $\mathcal{D}_{t}=\{Z_{1}^{t},\dots,Z_{N_{t}}^{t}\}$ from $P_{t}$, number of folds $K$.
2:Randomly split source indices $N_{s}]$ random folds of equal size: $\mathcal{I}_{1},\dots,\mathcal{I}_{K}$.
3:for $k\in[K]$ do
4:  Produce ML estimate $\widehat{\mu}^{(-k)}$ using $\mathcal{D}_{s,k}^{c}:=\mathcal{D}_{s}\setminus\mathcal{D}_{s,k}$, where $\mathcal{D}_{s,k}:=(Z^{s}_{j}:j\in\mathcal{I}_{k})$.
5:  Produce ML estimate $\widehat{\alpha}^{(-k)}$ using $\mathcal{D}_{s,k}^{c}$ and $\mathcal{D}_{t}$.
6:  Construct $\widehat{\theta}_{k}$ per Equation (1) with $\widehat{\mu}:=\widehat{\mu}^{(-k)},\widehat{\alpha}:=\widehat{\alpha}^{(-k)}$, and samples $\mathcal{D}_{s,k}$ and $\mathcal{D}_{t}$.
7:  Construct $\widehat{\sigma}_{k}$ per Equation (2) with $\widehat{\mu}:=\widehat{\mu}^{(-k)},\widehat{\alpha}:=\widehat{\alpha}^{(-k)}$, $\widehat{\gamma}:=\frac{N_{t}}{N_{s}}$, and samples $\mathcal{D}_{s,k}$ and $\mathcal{D}_{t}$.
8:Compute the average of the $K$ estimates: $\widehat{\theta}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\theta}_{k}$ and $\widehat{\sigma}^{2}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\sigma}_{k}^{2}$.
9:Return: Mean estimate $\widehat{\theta}$ and variance estimate $\widehat{\sigma}^{2}$.
Theorem 3.1.
Assume the learner has access to samples $Z_{1}^{s},\dots,Z_{N_{s}}^{s}\sim P_{s}$ and $Z_{1}^{t},\dots,Z_{N_{t}}^{t}\sim P_{t}$ satisfying Assumptions 12 and the assumptions of Theorems B.2 and B.4, all outlined in Appendix B. Then, letting $\widehat{\theta}$ and $\widehat{\sigma}^{2}$ denote the mean and variance returned by Algorithm 1, we have
$\widehat{\sigma}^{-1}\sqrt{N_{t}}\left(\widehat{\theta}-\theta\right)\Rightarrow\mathcal{N}(0,1).$
In particular, this implies that, for any $\delta\in(0,1)$, the set
$C_{1-\delta}:=\left[\widehat{\theta}-\frac{\widehat{\sigma}}{\sqrt{N_{t}}}z_{\delta/2},\widehat{\theta}+\frac{\widehat{\sigma}}{\sqrt{N_{t}}}z_{\delta/2}\right]$
is a $1-\delta$ confidence interval for $\theta$, where $z_{\delta}$ denotes the $\delta$ quantile of a standard normal R.V.
A complete theorem statement can be found in Theorem B.4. In Appendix C, we present a generalization to M-estimators (Theorems C.1 and C.4) and a corresponding proof — the above follows as a special case. We also provide examples of other target parameters, including rating variance and quantiles in Remark C.2 of the same appendix, which may be of independent interest.
The validity of Theorem 3.1 relies on several key assumptions, formally outlined in Appendix B. Of these, the most important is double robustness, which requires the product of nuisance estimation errors to decay at parametric (i.e. $\sqrt{N_{t}}$) rates. Formally, this assumption can be expressed via the condition that
$\|\widehat{\alpha}^{(-k)}-\alpha_{0}\|_{L^{2}}\cdot\|\widehat{\mu}^{(-k)}-\mu_{0}\|_{L^{2}}=o_{\mathbb{P}}(N_{t}^{-1/2})$ (3)
on each fold. See Appendix B for definition of $L^{2}$ norms and a formal definition of $o_{\mathbb{P}}$ notation (here, $o_{\mathbb{P}}(N_{t}^{-\beta})$ denotes convergence in probability at $N_{t}^{-\beta}$ rates). Notably, this condition allows each individual nuisance estimate to converge at non-parametric rates, thus permitting coverage even when estimates of either $\alpha_{0}$ or $\mu_{0}$ are of lower quality. For instance, one could have
$\max\left\{\|\widehat{\mu}^{(-k)}-\mu_{0}\|_{L^{2}},\|\widehat{\alpha}^{(-k)}-\alpha_{0}\|_{L^{2}}\right\}=o_{\mathbb{P}}(N_{t}^{-1/4})$
and still maintain valid coverage. We also note that the above convergence rate does not directly depend on the quality of persona ratings. Rather, the persona ratings serve as an extra covariate onto which we can regress human ratings $Y$. When persona ratings are highly correlated with human ratings, we may obtain faster convergence rates for $\hat{\mu}$. However, this does not prohibit convergence even when the quality of persona ratings is low. This phenomenon is illustrated in our experiments below.
Estimation Details. In Theorem 3.1 above, estimating $\mu_{0}$ is a standard regression task that can be accomplished using any off-the-shelf model (e.g. gradient boosted trees, neural networks). However, estimation of $\alpha_{0}(w,c)$, which is a complicated ratio of likelihood ratios and propensity scores, is more subtle. The standard approach for doubly-robust estimation would involve learning $\widehat{\omega}$ and $\widehat{\pi}$ separately (e.g., via gradient boosted trees) then estimating $\widehat{\alpha}$ by taking the ratio of predictions produced by each model. However, because $w$ can occupy a high dimensional (e.g., text) space, the variance in this ratio can be quite high. This variance in turn propagates to downstream estimates.
To address this challenge, we leverage a “Riesz loss” (Chernozhukov et al., 2022b; a; 2023) to estimate $\alpha_{0}$. Rather than learning $\omega_{0}$ and $\pi_{0}$ independently, the Riesz loss directly learns the ratio $\alpha_{0}(w,c)$. For our setting, letting $\beta_{0}(w):=\omega_{0}(w)/\pi_{0}(w)$, the Riesz loss minimizer is given by:
$\beta_{0}=\arg\min_{\beta}\left\{\mathbb{E}_{s}[C\cdot\beta(W^{s})^{2}]-2\mathbb{E}_{t}[\beta(W^{t})]\right\}.$ (4)
Therefore, to estimate $\alpha_{0}$, we minimize the finite-sample analogue of Equation 4 using $\mathcal{D}_{s,k}^{c}$ and $\mathcal{D}_{t}$ then plug this into Algorithm 1 (see Appendix D for details). As we show in $\S$ 4, this Riesz loss approach significantly improves the quality of downstream estimates.
Generalization to M-Estimators. Lastly, we note that our theoretical results also extend to target parameters $\theta_{t}$ defined as a solution to an M-estimation problem, i.e.
$0=\mathbb{E}_{t}[m(W,Y;\theta_{t})].$ (5)
For instance, if $m(w,y;\theta):=Q-\mathbbm{1}\{y\leq\theta\}$, then the solution $\theta_{t}$ gives the $Q$th quantile of $Y$ under $P_{t}$. In our running example, this corresponds to the 90’th quantile of “factual consistency” ratings provided over samples from the target distribution. Likewise, if
$m(w,y;(\rho,\theta)):=\begin{pmatrix}y-\rho\\ (y-\rho)^{2}-\theta\end{pmatrix}$,
then the solution $\theta_{t}$ to Equation (5) is $\rho_{t}=\mathbb{E}_{t}[Y]$ and $\theta_{t}=\mathrm{Var}_{t}[Y]$. This yields the variance in “factual consistency” ratings over $P_{t}$. Even without consideration of selection-bias, this result for M-estimation serves as a significant generalization of Chernozhukov et al. (2023), which only holds for regression problems. We describe our full approach for handling M-estimators, including assumptions, a cross-fitting algorithm, and nuisances, in Appendix C.
4 Experiments
[…]
  • Synthetic: All nuisance functions and target parameters are fully known (see Appendix E).
  • Semi-Synthetic: We sample 1000 real user conversations from PRISM (Kirk et al., 2024) and obtain the “ground truth” target parameter $\theta_{t}$ by treating ratings returned by an LLM-as-a-judge as human ratings ($Y$). We sample 50 such LLM ratings per item. Here, true nuisance functions are unknown. We sample persona ratings $\widehat{Y}$ by adding controlled error to the LLM-as-a-judge ratings (see $\S$ 4.1). This dataset instructs raters to assess the “helpfulness” of outputs on a 1-100 scale.
  • Real-World: We sample 300 real user conversations, rater characteristics (e.g., age, race), and human ratings ($Y$) from DICES (Aroyo et al., 2023), resulting in 300 conversations each with 25 human ratings each. We then sample persona ratings $\widehat{Y}$ by (i) adding controlled error to human ratings (see $\S$ 4.1) and (ii) via an LLM-as-a-judge with persona-based prompting. This dataset instructs raters to assess the “harmfulness” of outputs on a 1-4 scale.
In addition to providing a foundation for validating our doubly-robust estimator, the PSF offers a resource for the community to test future evaluation approaches under evaluation sampling bias.
4.1 Dataset Generation Procedure
We now describe how semi-synthetic (PRISM) and real-world (DICES) datasets are generated in the PSF. Further setup details, including prompts used for synthetic dataset generation, are reported in Appendix E.
Source and Target Populations. In the semi-synthetic dataset (PRISM), the source population consists of conversations where users are prompted to engage in controversial topics, while the target population consists of conversations with no guided prompts. In the real-world dataset (DICES), the source population contains 350 single-turn conversations flagged by safety experts as containing a single harm type (e.g., misinformation, legal), while the target contains more complex conversations rated as containing multiple types of harm. In both cases, we model each sample as a single user–system exchange extracted from a multi-turn dialogue. We embed the input–output pair from each exchange into a low-dimensional space by first applying an embedding model (MiniLM-L6-v2) then projecting to 15 dimensions via UMAP (Becht et al., 2019). [^4] We selected 15 dimensions to ensure embeddings retained some predictive signal for ratings and source/target membership while keeping dimensionality low; results remained stable for $\geq 12$ dimensions. We also vary the demographic composition of raters across populations. In both PRISM and DICES, we define the source distribution $P_{s}(X)$ using marginal probabilities of rater characteristics reported in DICES, and target distribution $P_{t}(X)$ using population statistics released in the U.S. Census Bureau’s 2022 Annual Social and Economic Supplement (Guzman & Kollar, 2023).
Covariate Shift. To vary the magnitude of covariate shift, we control the mixture between the source and target populations. We vary the content characteristics by controlling the proportion $\zeta\in[0,1]$ of target items contained within the source sample (sub-sampling from the full data to ensure that source and target sample sizes remain fixed). Additionally, we vary the rater distributions by taking the convex combination between all groups in each demographic stratum with normalization. The magnitude of the resulting covariate shift between samples is then given by the Sinkhorn Distance $\Delta\!\left(W^{s},W^{t}\right)$ (Feydy et al., 2019), where recall that $W=(X,V)$ and $V$ denotes the embedded content characteristics (MiniLM-L6-v2 + UMAP). We report the Sinkhorn distance normalized by subtracting the baseline case where there is no covariate shift for semi-synthetic and real-world settings, as it is inevitable that there will be variation in text embeddings despite sampling i.i.d. from pre-defined categories (e.g., harm types) within a population. This measure captures covariate shift resulting from content characteristics and demographic attributes in parallel.
Selection Bias. We model rater attrition—when raters fail to provide a rating due to failed attention checks or task abandonment—by varying the probability that each item is rated. In PRISM, we prompt the LLM to output both a rating and a non-response “refusal” flag. In DICES, we use rater self-assessments of task understanding to assign attrition scores (see Appendix E). We then transform attrition scores into dropout probabilities using a Beta CDF with $\alpha=3$ (increasing $\beta$ increases selection bias). We censor ratings according to these probabilities while retaining the “true” rating. We quantify the magnitude of selection bias via the dropout rate, i.e., the probability that a rater fails to rate an item. Crucially, the dropout rates we simulate mirror those observed in practice. In DICES, 19 of 123 raters (15.4%) were excluded due to failed attention checks, while in PRISM, 104 of 1500 raters (6.9%) failed to provide ratings after completing the background survey. As we show in our results, existing methods (e.g., RePPI (Ji et al., 2025)) exhibit severe miscoverage at these empirically observed dropout rates. This underscores the importance of correcting for selection bias.
Persona Quality. To systematically manipulate the quality of persona ratings, we perturb human ratings with controlled error. This error perturbation has (i) a bias parameter $\eta\in[-1,1]$, which induces a systematic shift, and (ii) a correlation parameter $\rho\in[-1,1]$, which parametrizes the Pearson correlation with human ratings. We construct the persona rating via a Cholesky-based procedure that transforms independent Gaussian noise to achieve the target correlation $\rho$:
$\hat{Y}=\text{clip}(\rho\cdot Y+\sqrt{1-\rho^{2}}\cdot Z\sigma_{Y}+\eta(y_{\max}-y_{\min}),y_{\min},y_{\max}),$
where $y_{\min},y_{\max}$ define the interval of the rating scale, and $Z$ represents independent Gaussian noise. To verify that our findings are robust to any artifacts of this perturbation process, we complement this error perturbation approach with persona ratings sampled from real LLMs.
4.2 Setup Details
Models. We use GPT-5 to generate synthetic “human” ratings for PRISM – i.e., used as $Y$ in our framework to obtain the ground truth target parameter $\theta_{t}$. We use Claude-{Haiku 3.5, Sonnet 3.5} and GPT-{5, 4o-Mini} to generate persona ratings for DICES. We report prompts, sampling temperature and decoding methods used for each LLM in Appendix E.
[…]
Metrics. We evaluate estimator quality using three metrics: Bias (MAE): $|\theta_{t}-\hat{\theta}_{t}|$, absolute deviation from the target parameter; Coverage: $\Pr(\theta_{t}\in[\hat{\theta}_{\mathrm{low}},\hat{\theta}_{\mathrm{high}}])$, the probability that the confidence interval covers the true parameter; and Interval Width: $\hat{\theta}_{\mathrm{high}}-\hat{\theta}_{\mathrm{low}}$, the length of the interval.
4.3 Results
Finding 1: DR (Riesz) obtains lower bias and improved coverage than baseline estimators. Figures 3 and 4 present our main findings varying (i) covariate shift, (ii) selection bias, and (iii) persona quality over all three datasets ($N=40$ trials per setting). We present cross-sectional results in Fig. 3 and average in Fig. 4. As illustrated in Fig. 3, DR (Riesz) obtains valid 95% CIs when (i) persona quality is high (top), (ii) covariate shift is moderate (middle) and (iii) across ranges of selection bias (bottom). In contrast, baseline estimators obtain valid coverage only on Synthetic when: (i) persona quality is very high (Fig. 3, top left) and (ii) and dropout rate is high (Fig. 3, bottom left). While counterintuitive, the second observation highlights the importance of examining covariate shift and selection bias in parallel; as dropout rate increases in Synthetic, the mean of remaining source samples more closely resembles that of the target distribution, leading to coincidentally higher coverage. Yet coverage remains poor on both PRISM and DICES.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Coverage by persona quality (top), covariate shift (center), and selection bias (bottom). DR (Riesz) attains better coverage than all baselines. Baselines with 0% coverage omitted to reduce clutter. $\eta=0.1$; $\rho=0.6$ for bottom two rows. Fig. 1012 (Appendix E) presents analogous results for Bias (MAE) and Interval Width.
Method Synthetic PRISM DICES
Bias Coverage Width Bias Coverage Width Bias Coverage Width
Sample Average 0.73 $\pm$ 0.17 0.06 $\pm$ 0.03 0.35 $\pm$ 0.01 1.30 $\pm$ 0.02 0.00 1.49 0.10 0.00 0.07
IPW 0.48 $\pm$ 0.05 0.00 1.00 $\pm$ 0.20 25.49 $\pm$ 0.02 0.00 2.62 $\pm$ 0.03 0.17 0.07 0.10
PAR 0.06 0.44 $\pm$ 0.03 0.10 0.83 $\pm$ 0.01 0.02 $\pm$ 0.01 0.91 0.05 0.04 0.02
Persona-Based 0.37 $\pm$ 0.01 0.00 0.17 10.00 $\pm$ 0.01 0.00 1.33 0.34 0.00 0.05
PPI++ 0.69 $\pm$ 0.16 0.03 $\pm$ 0.02 0.17 $\pm$ 0.01 1.03 $\pm$ 0.01 0.00 1.01 0.06 0.01 0.03
RePPI 0.10 $\pm$ 0.01 0.56 $\pm$ 0.09 0.19 0.63 $\pm$ 0.01 0.66 $\pm$ 0.02 1.36 0.04 $\pm$ 0.01 0.40 0.05
DR (Classical) 0.07 0.85 $\pm$ 0.01 0.21 0.68 $\pm$ 0.01 0.82 $\pm$ 0.02 1.75 0.05 0.32 0.06 $\pm$ 0.01
DR (Riesz) 0.03 1.00 0.28 $\pm$ 0.01 0.46 $\pm$ 0.01 0.93 $\pm$ 0.01 1.68 0.02 0.86 $\pm$ 0.01 0.09
Figure 4: Average Bias (MAE), Coverage, and Interval Width across experimental conditions presented in Fig. 3. Values in parentheses denote standard error (values <0.01<0.01 omitted to reduce clutter).
Finding 2: DR (Riesz) yields improved coverage and lower bias (MAE) than DR (Classical). Across levels of covariate shift, selection bias, and persona quality, we observe DR (Riesz) (Fig. 3; solid lines) obtains improved estimates compared to DR (Classical) (Fig. 3; dashed lines). While this behavior also appears in Synthetic (Fig. 3; left column), the gap between DR (Classical) and DR (Riesz) is especially pronounced when learning nuisance functions on embeddings of high-dimensional text (Fig. 3; PRISM, DICES). This illustrates the importance of directly estimating the re-weighting term $\alpha_{0}(W,C)$ (Equation. (4)) rather than learning $\omega_{0}(W)$ and $\pi_{0}(W)$ separately.
Finding 3: DR (Riesz) makes better use of persona ratings than baseline estimators. Several of our baselines — RePPI, PAR, and Persona-Based — use persona ratings to compute estimates. However, across levels of persona quality (Fig 3; top row), DR (Riesz) produces higher quality estimates than these baselines (with valid coverage for $\rho\geq.65$ on both PRISM and DICES). Fig. 5 extends this analysis to persona ratings obtained from real LLMs on DICES. For all LLMs, we observe that coverage of DR (Riesz) (solid) is markedly higher than that of RePPI (dashed). Further substantiating our systematic perturbation study Fig. 3 (top row), we observe that real LLMs that exhibit a higher correlation with human ratings (e.g., GPT-5; $\rho=0.43$) yield improved coverage over those with lower correlation (e.g., GPT-4o Mini; $\rho=0.23$). Furthermore, despite having lower correlation coefficients, we observe several models achieving comparable coverage to our artificially perturbed persona ratings (Fig. 3; $\rho=0.6$). Taken together, these findings illustrate that persona ratings from real LLMs-as-judges can be used to improve downstream estimates under evaluation sampling bias.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 5: Coverage of DR (Riesz) (solid) versus RePPI (dashed) when varying covariate shift (left) and selection bias (right) with persona ratings from different LLM judges. Parentheses denote Pearson correlation between persona and human ratings.
5 Related Work
We now provide a brief overview of related literature (see Appendix A for a detailed discussion).
Automated Evaluation with Persona Prompting. To address evaluation sampling bias, one strategy is to use an automated rater to rate outputs from the target distribution. Under this LLM-as-a-judge approach, a judge GenAI system rates the outputs of a target GenAI system (Li et al., 2024; Elangovan et al., 2024; Ye et al., 2024; Bubeck et al., 2023; Zheng et al., 2023). Because human raters often disagree on criteria such as “helpfulness” or “relevance” (Kirk et al., 2024), prior work has explored instructing judge systems to adopt personas — descriptions of humans with specific characteristics (Castricato et al., 2024; Fröhling et al., 2024; Orlikowski et al., 2025; Deng et al., 2025). However, work has also shown that persona ratings are often an imperfect proxy for human ratings (Santurkar et al., 2023; Neumann et al., 2025). Thus, our work treats persona ratings as a useful yet incomplete proxy for human raters to improve GenAI system quality estimates.
Frameworks for Sample Efficient Estimation. Other works propose methods for improving statistical inference when data is scarce but ML predictions are abundant. Prediction-Powered Inference (PPI) and its computationally efficient variant PPI++ use ML predictions to tighten confidence intervals through a “rectifier term” that corrects for bias in ML predictions (Angelopoulos et al., 2023a; Chatzi et al., 2024; Fisch et al., 2024; Angelopoulos et al., 2023b). Ji et al. (2025) show PPI++ to be a special case of M-estimation with surrogate outcomes, a classical problem in causal inference (Robins et al., 1994; Robins & Rotnitzky, 1995; Tsiatis, 2006), and in turn propose recalibrated PPI (or RePPI) to offer more efficient estimation. However, these approaches fail to give valid coverage under evaluation sampling bias. We develop a doubly-robust estimator (Bang & Robins, 2005; Chernozhukov et al., 2018; 2023) that can handle covariate shift and selection bias simultaneously while making use of surrogate predictions/persona ratings. We also use “Riesz losses” (Chernozhukov et al., 2023; 2022a; 2022b) to estimate complicated nuisance parameters using generic ML learners.
6 Conclusion
Our work answers calls for greater consideration of external validity concerns in Generative AI evaluation (Weidinger et al., 2025; Ibrahim et al., 2024; Liao et al., 2021; Salaudeen et al., 2025) through a theoretically rigorous and empirically validated estimation framework. Our framework provides a path forward for combining limited human ratings observed under sampling bias with imperfect persona ratings to obtain statistically valid system quality estimates. Beyond our specific doubly-robust estimation framework, our Persona Simulation Framework (PSF) also provides a reusable community resource for validating future methods designed to address sampling bias. While our framework relaxes the MCAR assumption imposed by existing estimation frameworks, it also imposes assumptions — e.g., no concept drift — on the evaluation process. Future work should also consider how violations of this and other assumptions in Appendices B and C might affect system quality estimates.
7 Reproducibility Statement
We take several steps to ensure the reproducibility of our work. First, we document all theoretical assumptions required by our framework and provide complete proofs in Appendix B. Second, we provide thorough documentation of our experiment design, hyperparameters, and datasets required to reproduce our empirical results in Appendix E. Finally, we provide code necessary to reproduce our analysis in the supplementary material. We plan to release all code and datasets in our Persona Simulation Framework along with the paper so that the broader community can build upon our framework.
8 Contribution Statement
KT: Led all work on experiments. Designed and implemented the Persona Simulation Framework, collected synthetic datasets from LLMs, and implemented estimators included in the experimental results. Supported drafting of the main text.
JW: Led all work on theoretical development and validation of the doubly-robust estimator. Stated and proved all results. Drafted appendices B, C, and D. Conducted early-stage experiments on synthetic data. Supported drafting of the main text.
LG: Led conceptual framing of the project. Designed the problem formulation. Supported implementation of the Persona Simulation Framework (e.g., data generation pipeline, Riesz loss). Wrote the main text with support from JW and KT.
Appendix A Extended Related Work
A.1 Threats to the External Validity of Generative AI Evaluations
In the quantitative social sciences, external validity describes the extent to which findings from a study generalize to different populations, settings, and times (Findley et al., 2021). Threats to external validity have long been studied in survey research. For example, Levay et al. (2016) found notable discrepancies between the demographic composition of convenience samples obtained via Amazon Mechanical Turk (MTurk) versus nationally representative American National Election Study (ANES) samples. While Mullinix et al. (2015) observe that study findings often remain robust to such discrepancies, (Zhou & Fishbach, 2016) demonstrate that differential non-compliance — a form of selection bias in which participants drop out from studies non-randomly — can have substantive effects on studies’ results. Myriad factors contribute to this selection bias, such as participants’ motivation and language skills (Goodman & Paolacci, 2017).
More recently, concerns have emerged surrounding the external validity of evaluations obtained from general purpose benchmarks (e.g., MMLU, BigBench) and leaderboards (e.g., Chatbot Arena) (Ibrahim et al., 2024; Ouyang et al., 2023; Liao & Xiao, 2023). As with survey research, GenAI performance measurements can be subject to covariate shift when the distribution of system outputs or human raters differs between a lab-based evaluation and target deployment context (Saad-Falcon et al., 2023; Leemann et al., 2024; Kirk et al., 2024). Likewise, differential non-compliance can occur when raters in online rating platforms drop-out due to failed quality checks (e.g., due to poor English language proficiency) (Hsueh et al., 2009). Such selection bias can confound results if common factors (e.g., English language proficiency) affect rater drop-out and their ratings. Selection bias can also arise if some raters are more likely to voluntarily assign ratings than others — e.g., when busy physicians rate complex system outputs less frequently than more available medical students. While growing work has highlighted external validity as an important desideratum for evaluations (Ibrahim et al., 2024; Ouyang et al., 2023; Liao & Xiao, 2023), to our knowledge, no existing statistical frameworks simultaneously address covariate shift, selection bias, and high-dimensional model outputs while leveraging imperfect automated ratings.
We address this gap by developing a statistical framework for characterizing and mitigating threats to the external validity of GenAI performance evaluations. We devise a data-efficient estimator that corrects for covariate shift and selection bias in parallel, given ratings from a source population and predictions generated by a black-box machine learning model over both source and target populations. Some advances in our framework provide a new perspective on classic methodological challenges in survey research. For example, we provide a doubly-robust alternative to the reweighting estimators traditionally used to correct for selection bias. This approach obtains valid coverage even when the re-weighting model is misspecified. Our framework also addresses novel challenges that arise in the GenAI evaluation context. In particular, we leverage embeddings to support robust statistical inference over high-dimensional model output spaces (e.g., text, image) as opposed to the structured data formats traditionally used for survey research. Central to our approach is the principled adoption of synthetic ratings generated by an AI persona, which we discuss next.
A.2 Automated Evaluation with LLM-as-a-Judge and Persona Prompting
Given the cost and scalability challenges associated with collecting human ratings, automated methods are increasingly used to scale up evaluation workflows traditionally performed by humans. In particular, the LLM-as-a-judge paradigm introduces a second judge GenAI system to evaluate the outputs returned by a target GenAI system (Li et al., 2024; Elangovan et al., 2024; Ye et al., 2024; Bubeck et al., 2023; Zheng et al., 2023). Because human raters can disagree as to whether a model output is “helpful”, or “relevant” (Kirk et al., 2024), recent work has proposed instructing judge systems to adopt personas–that is, descriptions of humans with specific sociodemographic characteristics, such as gender and race (Castricato et al., 2024; Dong et al., 2024; Fröhling et al., 2024; Wright et al., 2024; Orlikowski et al., 2025; River Dong et al., 2025). This persona-based prompting strategy is designed to better-account for sources of rater-specific variation throughout the evaluation process.
These automated evaluation methods offer a promising approach to mitigate the external validity threats described in $\S$ A.1. In particular, judge systems with persona prompting can generate low-cost synthetic ratings when human ratings from the target population are limited. However, because such ratings may be systematically biased (Santurkar et al., 2023; Neumann et al., 2025), their direct adoption in evaluation pipelines may yield biased performance measurements. Our proposed doubly-robust approach addresses this challenge by treating LLM-as-a-judge ratings as potentially informative yet biased proxies for human ratings. This approach combines surrogate ratings with human ratings (observed under evaluation sampling bias) to obtain statistically valid confidence intervals in the target population of interest.
A.3 GenAI Systems as Human Surrogates in Social Science Studies
While our work foregrounds GenAI evaluation challenges, it bears conceptual and methodological similarity to work investigating GenAI systems as surrogates for human subjects in social science studies. Inline with the turn towards crowdworkers as a low-cost surrogate for target study populations ($\S$ A.1), this growing line of work introduces GenAI systems as a surrogate for more costly human subjects (Argyle et al., 2023). Notably, such work often targets the very same statistical parameters recovered by our general M-estimation framework (Table LABEL:tab:parameters). For example, let $v$ denote an item in an opinion poll (e.g., “do you believe in the right to bear arms?”), let $X$ denote rater characteristics (e.g., locale and demographics, per (Santurkar et al., 2023)), and let $Y$ represent a binary response (endorse/not endorse) to the survey item. The parameter
$\theta_{t}(v):=\mathbb{E}_{t}[Y\mid V=v]$ (6)
denotes the proportion of raters in the target population who endorse this survey item. Thus, researchers can also leverage our methodology when using GenAI systems as surrogates for human subjects in social science studies. Given a finite sample of human ratings from the source population $\{(X_{i},V_{i},Y_{i})\}_{i=1}^{N}\sim P_{s}$, and surrogate data produced for the source and target population, researchers can recover informative and statistically valid confidence intervals for parameters defined over the target population of human subjects.
Given the overlap between our motivating application and social science studies, we also discuss methods advancing the principled adoption of GenAI systems as surrogate data in social science research (Broska et al., 2024; Egami et al., 2024; 2023). These works view surrogate data as a flawed (Bisbee et al., 2024; Park et al., 2024; Takemoto, 2024; Abdurahman et al., 2024) but potentially informative source of information for statistical inference in social science studies. For instance, Broska et al. (2024) leverage prediction powered inference (Angelopoulos et al., 2023a) to correct for bias in surrogate data. However, as discussed in $\S$ A.4, PPI is vulnerable to covariate shift and selection bias between source and target study populations. Most related to our work, (Egami et al., 2024; 2023) introduce a doubly robust estimation approach that generalizes PPI by applying a bias-correction to the underlying moment function (as opposed to the outcome variable). Critically, however, this approach takes a design-based sampling procedure, which assumes that the probability of labeling a sample is known by the researcher in advance. This precludes the more general setting we study in our work, in which the reweighting function is unknown in advance. [^5] As noted in (Egami et al., 2024; 2023), this design-based sampling approach is well-motivated when the full corpus of documents to be annotated and corresponding sampling probabilities are known in advance. However, this assumption is violated in our setting, in which the true source/target distribution weights are unknown. As a result, the framework proposed by (Egami et al., 2024; 2023) is not directly applicable in our motivating setting with evaluation sampling bias. Having discussed both GenAI and social science applications of our framework, we now turn to the underlying statistical methodology we advance in this work.
A.4 Statistical Frameworks for Sample Efficient-Estimation
Recent developments in black-box predictive models that can operate on multi-modal representations has spurred significant interest in how these predictions might be used to improve statistical inference (Angelopoulos et al., 2023a; b; Fisch et al., 2024; Eyre & Madras, 2024; Dorner et al., 2024; Ji et al., 2025; Saad-Falcon et al., 2023; Fogliato et al., 2024). These frameworks address the challenge of making valid statistical inferences when labeled data is scarce but black-box predictions cheap and abundant. We briefly review developments in this literature before identifying key gaps addressed by our approach.
Prediction-Powered Inference (PPI) uses predictions from a black-box machine learning model to tighten confidence intervals when labeled data is scarce (Angelopoulos et al., 2023a; b). This is done through the addition of a “rectifier” — a mean zero term that contrasts the performance of the model’s predictions on the labeled and unlabeled points. While initial variants of PPI were not computationally efficient, Angelopoulos et al. (2023b) introduce a PPI++ framework, which introduces a “trust” parameter $\lambda$ to control the magnitude of the rectifier. This allows for efficiently computable confidence sets that are provably tighter than those computed just from labeled data. We also emphasize recent work due to Ji et al. (2025), which shows that PPI++ is just a special parametric class of solutions for M-estimation with surrogates outcomes — a classical, well-studied problem in the causal inference/missing data literature (Robins et al., 1994; Robins & Rotnitzky, 1995; Tsiatis, 2006). While the authors do not directly mention Neyman orthogonal scores (Neyman, 1979; Chernozhukov et al., 2018) in their work, they construct what is implicitly a Neyman orthogonal score for the problem at hand and propose a solution based on cross-fitting. Additional theoretical developments along these lines have been proposed — Ao et al. (2024) propose a framework for adaptive estimation of linear functionals based on supplied predictions. Likewise, Xu et al. (2025) consider a general semi-parametric framework for estimating functionals of the data generating distribution in the presence of ML predictions. However, none of the aforementioned works provides a framework usable in settings where (a) unlabeled and labeled samples come from different distributions (i.e. covariate shift) and (b) data is missing at random (MAR), (i.e. the probability that outcomes are observed for any given individual depend on their features). We close this gap by proposing a doubly-robust estimator (Bang & Robins, 2005) and a general algorithm based on cross-fitting (Chernozhukov et al., 2018) for solving M-estimation problems in the presence of both covariate shift and heterogeneity in data missingness. We also incorporate recent ideas on “Riesz losses” (Chernozhukov et al., 2022b; 2023; a; Hirshberg & Wager, 2021), loss functions that specify complicated nuisance functions as their minimizers. By using Riesz losses to learn the re-weighting function $\alpha_{0}(W,C)$, we avoid constructing plug-in estimates for $\omega_{0}$ and $\pi_{0}$ and computing their quotient, which can result in high bias.
Appendix B Assumptions, Notation, and Results From Section 3
B.1 Assumptions on Data-Generating Distributions
We start by formally describing the data generating processes considered throughout the paper. We start by describing the assumptions we place of the “full-data” source and target distributions.
Assumption 1.
We assume there are two “full data” distributions over tuples $(X,V,C,\widehat{Y},Y)$: a source distribution $P_{s}$ and a target distribution $P_{t}$. For simplicity, we let $W=(X,V)$ denote the extended set of covariates, and assume $W\in\mathcal{W}$ where $\mathcal{W}$ is some generic measurable space. We assume the following hold:
  1. 1.
    (No Concept Drift) The conditional distribution of $Y$ given $W$ is the same under $P_{s}$ and $P_{t}$, i.e. for any $w$
    $P_{s}(Y\in E\mid W=w)=P_{t}(Y\in E\mid W=w)$
    for any event $E$.
  2. 2.
    (Surrogates are Functions of the Data) The observed surrogate $\widehat{Y}$ satisfies:
    $\widehat{Y}=f(X,V,\epsilon),$
    where $\epsilon$ is a random variable independent of the vector $(X,V,C,Y)$.
  3. 3.
    (Positivity) We have
    $0<\pi_{0}(W)\leq 1\qquad\text{where}\qquad\pi_{0}(w):=P_{s}(C=1\mid W=w).$
  4. 4.
    (Conditional Ignorability Under Source) The outcome $Y$ is conditionally independent of $C$ given extended covariates $W$, i.e. we have
[…]
    where the conditional independence is under $P_{s}$.
  5. 5.
    (Overlap) The likelihood/density ratio of $W$ between $P_{t}$ and $P_{s}$, defined as
    $\omega_{0}(w):=\frac{dP_{t}}{dP_{s}}(w)$
    exists and is finite almost surely.
We may interchangeably write $P^{b}$ or $P_{b}$ for $b\in\{s,t\}$ as the distribution over source and target samples, and $\mathbb{E}^{b}$ and $\mathbb{E}_{b}$ interchangeably as the corresponding expectations. We typically write $(X^{b},V^{b},C^{b},Y^{b},\widehat{Y}^{b})$ and $W^{b}=(X^{b},V^{b})$ for samples drawn from $P_{b}$.
We briefly parse the above assumptions. The first simply says that even if the distribution of $W=(X,V)$ changes wildly, the conditional distribution of $Y$ remains the same. The second assumption regards the predictions/surrogate outcomes and is trivially satisfied if $\widehat{Y}$ is the prediction of a generative AI model that depends only on $W$ and independent, external sources of randomness. The third assumption, positivity, states that under the source distribution there is always some probability we observe the true outcome. The fourth assumption is an analogue to conditional ignorability from the causal inference literature, and says that the outcome $Y$ is conditionally independent of whether or not data is observed given covariates. We note that, under the above assumptions, $\widehat{Y}$ is also conditionally independent of $Y$ and $C$ given $W$. Lastly, the final assumption guarantees overlap, or that the support of $P_{t}$ is contained in the support of $P_{s}$ — a necessary assumption in order to perform inference under covariate shift.
The above assumption concerns fully-observed data — in practice, the learner will only ever observe outcomes for samples where $C=1$. That is, there is partial-observation of outcomes in the source population, but outcomes are entirely absent in the target population. We formalize this in the following assumption.
Assumption 2.
The learner only ever observes $Y$ for samples where $C=1$. In other words, observed samples form each distribution take the following form:
  1. 1.
    (Source Samples) The learner observes samples of the form $Z^{s}=(X^{s},V^{s},C^{s},C^{s}\cdot Y^{s},\widehat{Y}^{s})$ from $P_{s}$.
  2. 2.
    (Target Samples) The learner observes samples of the form $Z^{t}=(X^{t},V^{t},\widehat{Y}^{t})$ from $P_{t}$.
An important consequence of Assumption 1 is that, even when only observe partial data (per Assumption 2), we can still identify general classes of estimands under the target distribution $P_{t}$. This is clarified in the following lemma.
Lemma B.1.
Let $f$ be an arbitrary function of $(Y,\widehat{Y},W)$. Then, we have
$\mathbb{E}_{t}[f(Y^{t},\widehat{Y}^{t},W^{t})]=\mathbb{E}_{s}[\alpha_{0}(W^{s},C^{s})f(Y^{s},\widehat{Y}^{s},W^{s})],$
where $\alpha_{0}(w,c):=c\frac{\omega_{0}(w)}{\pi_{0}(w)}$.
Proof.
Observe that we have:
$\mathbb{E}_{s}[\alpha_{0}(W^{s},C^{s})f(Y^{s},\widehat{Y}^{s},W^{s})]$ $=\mathbb{E}_{s}\left[C^{s}\frac{\omega_{0}(W^{s})}{\pi_{0}(W^{s})}f(Y^{s},\widehat{Y}^{s},W^{s})\right]$
$=\mathbb{E}_{s}\left[\frac{\omega_{0}(W^{s})}{\pi_{0}(W^{s})}\mathbb{E}\left(C^{s}f(Y^{s},\widehat{Y}^{s},W^{s})\mid W^{s}\right)\right]$
$=\mathbb{E}_{s}\left[\frac{\omega_{0}(W^{s})}{\pi_{0}(W^{s})}\mathbb{E}\left(C^{s}\mid W^{s}\right)\mathbb{E}\left(f(Y^{s},\widehat{Y}^{s},W^{s})\mid W^{s}\right)\right]$
$=\mathbb{E}_{s}\left[\omega_{0}(W^{s})\mathbb{E}\left(f(Y^{s},\widehat{Y}^{s},W^{s})\mid W^{s}\right)\right]$
$=\mathbb{E}_{s}\left[\omega_{0}(W^{s})f(Y^{s},\widehat{Y}^{s},W^{s})\right]$
$=\mathbb{E}_{t}[f(Y^{t},\widehat{Y}^{t},W^{t})].$
[…]
B.2 Notation
We now discuss some additional notation that will be leveraged in the sequel.
We will need to condition on independent, random nuisance estimates regularly in the sequel. For $b\in\{s,t\}$, if $U$ is another random variable (e.g. $U=\widehat{g}$ where $\widehat{g}$ denotes a generic nuisance estimate) and $f(Z,U)$ is some generic function, we define $P^{b}_{Z}$ and $\mathbb{E}^{b}_{Z}$ as the distribution and expectation over just the randomness in $Z$ while conditioning on $U$, i.e.
$P^{b}_{Z}(f(Z,U)\in E):=P_{b}(f(Z,U)\in E\mid U)\quad\text{and}\quad\mathbb{E}^{b}_{Z}f(Z,U):=\mathbb{E}^{b}_{Z}(f(Z,U)\mid U).$
We define the empirical distributions with respect to observations as $\mathbb{P}_{N_{s}}:=\frac{1}{N_{s}}\sum_{j=1}^{N_{s}}\delta_{Z_{i}^{s}}$ and $\mathbb{P}_{N_{t}}:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\delta_{Z_{i}^{t}}$, where $\delta_{z}$ denotes the point-mass distribution on $z$. Thus, for a general random function $\widehat{g}$ of data $Z^{b}$, we have $\mathbb{P}_{N_{s}}\widehat{g}(Z^{s}):=\frac{1}{N_{s}}\sum_{j=1}^{N_{s}}\widehat{g}(Z_{j}^{s})$ and $\mathbb{P}_{N_{t}}\widehat{g}(Z^{t}):=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{g}(Z_{i}^{t})$. We define the $L^{2}(P_{Z}^{b})$ norm of a potentially random $\mathbb{R}^{d}$-valued function $\widehat{g}$ depending on a subset of features $S\subset Z$ as
$\|\widehat{g}\|_{L^{2}(P^{b}_{Z})}:=\left(\mathbb{E}_{Z}^{b}\left[\|\widehat{g}(S)\|_{2}^{2}\right]\right)^{1/2}=\left(\mathbb{E}_{b}\left(\|\widehat{g}(S)\|_{2}^{2}\mid\widehat{g}\right)\right)^{1/2}.$
We likewise define the $L^{\infty}(P_{Z}^{b})$ norm $\widehat{g}$ as the $\|\widehat{g}\|_{L^{\infty}(P^{b}_{Z})}:=\inf\left\{b\in\mathbb{R}:P^{b}_{Z}(\|\widehat{g}\|_{\infty}>b)=0\right\}$, where $\|x\|_{\infty}:=\max\{x_{1},\dots,x_{d}\}$. Note that whenever $\widehat{g}$ is random, these norms are random variables as well.
Given a (random or deterministic) sequence $(X_{n})_{n\geq 1}$ in a normed space $(\mathcal{X},\|\cdot\|)$ and a deterministic scalar sequence $(b_{n})_{n\geq 1}$, we say $X_{n}=o(b_{n})$ if $\lim_{n\rightarrow\infty}\frac{\|X_{n}\|}{b_{n}}=0$ almost surely and $X_{n}=O(b_{n})$ if there exists a constant $B>0$ such that $\frac{\|X_{n}\|}{b_{n}}\leq B$ for all $n\geq 1$. We say a sequence of random variables $(X_{n})_{n\geq 1}$ converges in probability to zero, denoted by $X_{n}\xrightarrow[n\rightarrow\infty]{\mathbb{P}}0$, if we have
$\lim_{n\rightarrow\infty}\mathbb{P}(\|X_{n}\|\geq\epsilon)=0\qquad\text{for any }\epsilon>0.$
We say $X_{n}=o_{\mathbb{P}}(b_{n})$ if $X_{n}/b_{n}\xrightarrow{\mathbb{P}}0$, and $X_{n}=O_{\mathbb{P}}(b_{n})$ if for any $\epsilon>0$, there is a constant $M_{\epsilon}>0$ such that $\limsup_{n\rightarrow\infty}\mathbb{P}(\|X_{n}\|/b_{n}\geq M_{\epsilon})\leq\epsilon$.
If $(X_{n})_{n\geq 0}$ is a sequence of random variables in $\mathbb{R}^{d}$, we always refer convergence in probability with respect to the $\ell_{2}$-norm, where $\|x\|_{p}:=\left(\sum_{k=1}^{d}x_{k}^{p}\right)^{1/p}$ for any $1\leq p<\infty$. Likewise, if $(X_{n})_{n\geq 0}$ is a sequence of random matrices, we assume convergence in probability is defined with respect to the operator norm $\|X\|_{op}:=\sup_{\begin{subarray}{c}u\in\mathbb{R}^{d}\\ \|u\|_{2}=1\end{subarray}}\|Xu\|_{2}$. For a non-singular matrix $A\in\mathbb{R}^{d\times d}$, we let $A^{-1}$ denote its inverse, and $A^{-T}$ denote the transpose of the inverse. If $x\in\mathbb{R}^{d}$ is a vector, we let $x^{\otimes 2}:=xx^{\top}$ for convenience.
B.3 General De-Biased Inference for an Expected Outcome
We now state and prove our main theorem for performing inference on an expected outcome $\theta_{t}=\mathbb{E}_{t}[Y^{t}]$ under evaluation sampling bias (i.e. covariate shift and selection bias). We start by stating a result that assumes the learner is given nuisance estimates that are independent of the entire sample of data. In the sequel, we describe an extended, cross-fitting based result that makes more efficient use of the data.
Theorem B.2.
Suppose Assumption 1 holds, and assume the learner has access to mutually independent samples $Z_{1}^{s},\dots,Z_{N_{s}}^{s}$ from $P_{s}$ and $Z_{1}^{t},\dots,Z_{N_{t}}^{t}$ from $P_{t}$, as outlined in Assumption 2. Let $\mu_{0}(w)$ and $\alpha_{0}(w,s)$ be true, unknown nuisances given by
$\mu_{0}(w):=\mathbb{E}_{t}[Y^{t}\mid W^{t}=w]=\mathbb{E}_{s}[Y^{s}\mid W^{s}=w]\quad\text{and}\quad\alpha_{0}(w,c)=\frac{c\omega_{0}(w)}{\pi_{0}(w)},$
where $\pi_{0}(w):=P_{s}(C^{s}=1\mid W^{s}=w)$ and $\omega_{0}(w)=\frac{dP_{t}}{dP_{s}}(w)$. Assume the following conditions hold.
  1. 1.
    (Ratio of Sample Sizes) There is some constant $0<\gamma<\infty$ such that $N_{t}/N_{s}\rightarrow\gamma$.
  2. 2.
    (Nuisance Convergence) We have access to estimates $\widehat{\mu},\widehat{\alpha}$ that are independent of the sample such that
    $\|\widehat{\mu}-\mu_{0}\|_{L^{2}(P_{Z}^{b})},\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{b}_{Z})}=o_{\mathbb{P}}(1)$
    and
    $\|\widehat{\mu}-\mu_{0}\|_{L^{2}(P^{b}_{Z})}\cdot\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{b}_{Z})}=o_{\mathbb{P}}(N_{t}^{-1/2})$
    for $b\in\{s,t\}$. Further, we assume $\widehat{\alpha}(w,0)=0$.
  3. 3.
    (Boundedness) The representer $\alpha_{0}(W^{s},C^{s})$ and the outcomes $Y^{s}$ are almost surely bounded.
Let the estimator $\widehat{\theta}$ be defined via the de-biased equation
$\widehat{\theta}:=\mathbb{P}_{N_{t}}\widehat{\mu}(W^{t},\widehat{Y}^{t})+\mathbb{P}_{N_{s}}\widehat{\alpha}(W^{s},C^{s})\left\{Y^{s}-\widehat{\mu}(W^{s},\widehat{Y}^{s})\right\}.$
Then, we have asymptotic linearity, i.e.
$\sqrt{N_{t}}(\widehat{\theta}-\theta_{t})=\frac{1}{\sqrt{N_{t}}}\sum_{i=1}^{N_{t}}\mu_{0}(W^{t}_{i})+\frac{\gamma}{\sqrt{N_{t}}}\sum_{j=1}^{N_{s}}\alpha_{0}(W^{s}_{j},C^{s}_{j})\left\{Y^{s}_{j}-\mu_{0}(W^{s}_{j})\right\}+o_{\mathbb{P}}(1).$
Furthermore, we have
$\sqrt{N_{t}}(\widehat{\theta}-\theta_{t})\Rightarrow\mathcal{N}(0,\sigma^{2}),$
so long as the asymptotic variance, given by
$\sigma^{2}=\mathrm{Var}_{t}[\mu_{0}(W^{t})]+\gamma\mathbb{E}\left[\alpha_{0}(W^{s},C^{s})^{2}\mathrm{Var}_{s}[Y^{s}\mid W^{s}]\right],$
is non-zero.
The following corollary shows that one can use the plug-in variance estimate to construct asymptotically valid confidence intervals.
Corollary B.3.
Under the same assumptions of Theorem B.2, the plug-in variance estimate
$\widehat{\sigma}^{2}:=\mathbb{P}_{N_{t}}\left\{\widehat{\mu}(W^{t},\widehat{Y}^{t})-\overline{\mu}\right\}^{2}+\frac{N_{t}}{N_{s}}\mathbb{P}_{N_{s}}\widehat{\alpha}(W^{s},C^{s})^{2}\left\{Y^{s}-\widehat{\mu}(W^{s})\right\}^{2}$
is consistent, where $\overline{\mu}:=\sum_{i=1}^{N_{t}}\widehat{\mu}(W^{t},\widehat{Y}^{t})$. Consequently, if the asymptotic variance $\sigma^{2}$ is non-zero, we have
$\frac{\sqrt{N_{t}}}{\widehat{\sigma}}(\widehat{\theta}-\theta_{t})\Rightarrow\mathcal{N}(0,1),$
and thus
$C_{1-\delta}:=\left[\widehat{\theta}-\frac{\widehat{\sigma}}{\sqrt{N_{t}}}z_{\delta/2},\widehat{\theta}+\frac{\widehat{\sigma}}{\sqrt{N_{t}}}z_{\delta/2}\right]$
is a $1-\delta$ confidence interval for $\theta_{t}$, where $z_{\delta}$ denotes the $\delta$ quantile of a standard normal random variable.
The proofs of Theorem B.2 and Corollary B.3 follow immediately from applying Theorem C.1 and Corollary C.3 in Appendix C (which concerns the case of general M-estimation) to the score $m(w,y;\theta):=y-\theta$.
B.4 Cross-Fitting for Means
We now provide a cross-fitting algorithm for estimating $\theta_{t}=\mathbb{E}_{t}[Y]$ and state an analogue of Theorem B.2. In short, cross-fitting works by splitting the data in $K$ folds of roughly equal size. If $\mathcal{I}_{k}$ denotes the $k$th fold of data, the algorithm uses all data outside the $k$th fold (so, in the complement set $\mathcal{I}_{k}^{c}$) to construct nuisance estimates. These nuisance estimates are then used to estimate the mean on the $k$th fold. This splitting strategy ensures that, on each fold, the nuisance estimates and transformed data are independent of one another. This allows one to apply the asymptotic linearity result of Theorem B.2 on each fold to asymptotic normality of the cross-fitting estimate.
We now state the cross-fitting algorithm (Algorithm 2) and corresponding convergence theorem (Theorem B.4). The proof of the cross-fitting result follows from Theorem C.4 in Appendix C, a generic result on cross-fitting for M-estimators under sampling bias.
Algorithm 2 Doubly-Robust Estimator with $K$-fold Cross-Fitting (detailed version of Algorithm 1)
1:Input: Samples $\mathcal{D}_{s}:=\{Z_{1}^{s},\dots,Z_{N_{s}}^{s}\}$ from $P_{s}$, samples $\mathcal{D}_{t}:=\{Z_{1}^{t},\dots,Z_{N_{t}}^{t}\}$ from $P_{t}$, number of folds $K$.
2:Randomly split source indices $N_{s}]$ into random folds of equal size: $\mathcal{I}_{1},\dots,\mathcal{I}_{K}$.
3:for $k\in[K]$ do
4:  Produce ML regression estimate $\widehat{\mu}^{(-k)}$ using $\mathcal{D}_{s,k}^{c}$, where $\mathcal{D}_{s,k}:=(Z^{s}_{i}:i\in\mathcal{I}_{k})$.
5:  Produce ML nuisance estimate $\widehat{\alpha}^{(-k)}$ using $\mathcal{D}_{s,k}^{c}$ and $\mathcal{D}_{t}$.
6:  Produce parameter and variance estimates:
$\widehat{\theta}_{k}$ $:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\mu}^{(-k)}(W_{i}^{t},\widehat{Y}_{i}^{t})$
$\quad+\frac{K}{N_{s}}\sum_{j\in\mathcal{I}_{k}}\widehat{\alpha}^{(-k)}(W_{j}^{s},C_{j}^{s})\left\{Y_{j}^{s}-\widehat{\mu}^{(-k)}(W_{j}^{s},\widehat{Y}_{j}^{s})\right\},$
$\widehat{\sigma}^{2}_{k}$ $:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\left\{\widehat{\mu}^{(-k)}(W_{i}^{t},\widehat{Y}_{i}^{t})-\widehat{m}^{t}_{k}\right\}^{2}$
$\quad+\frac{N_{t}}{N_{s}}\frac{K}{N_{s}}\sum_{j\in\mathcal{I}_{k}}\widehat{\alpha}^{(-k)}(W_{j}^{s},C_{j}^{s})^{2}\left\{Y_{j}^{s}-\widehat{\mu}^{(-k)}(W_{j}^{s},\widehat{Y}_{j}^{s})\right\}^{2},$
where $\widehat{m}^{t}_{k}:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\mu}^{(-k)}(W_{i}^{t},\widehat{Y}_{i}^{t})$.
7:Compute the average of the $K$ estimates: $\widehat{\theta}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\theta}_{k}$ and $\widehat{\sigma}^{2}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\sigma}_{k}^{2}$.
8:Return: Mean estimate $\widehat{\theta}$ and variance estimate $\widehat{\sigma}^{2}$.
Theorem B.4.
Assume the same setup as Theorem B.2, and let $\widehat{\mu}^{(-k)},\widehat{\alpha}^{(-k)}$, $\widehat{\theta}$, and $\widehat{\sigma}$ be as in Algorithm 2. Further, suppose the second assumption of Theorem B.2 holds for each nuisance estimate $\widehat{\mu}^{(-1)},\dots,\widehat{\mu}^{(-K)}$ and $\widehat{\alpha}^{(-1)},\dots,\widehat{\alpha}^{(-K)}$.Then, we have
$\frac{\sqrt{N_{t}}}{\widehat{\sigma}}(\widehat{\theta}-\theta_{t})\Rightarrow\mathcal{N}(0,1).$
Thus, the set $C_{1-\delta}$ defined in Corollary B.3 still serves as a $1-\delta$ asymptotic confidence interval for $\theta_{t}$.
Appendix C General M-estimators Under Covariate Shift
In this appendix, we prove a general result on the convergence of de-biased M-estimators under covariate and differential non-compliance. Our results from the previous appendix, which regarded the special case where the target parameter was the expected outcome $\theta_{t}=\mathbb{E}_{t}[Y]$, follow as a special case of the following.
Theorem C.1.
Suppose Assumption 1 holds, and assume the learner has access to mutually independent samples $Z_{1}^{s},\dots,Z_{N_{s}}^{s}$ from $P_{s}$ and $Z_{1}^{t},\dots,Z_{N_{t}}^{t}$ from $P_{t}$, as outlined in Assumption 2. Let $\psi_{0}(w)$ and $\alpha_{0}(w,c)$ be the true, unknown nuisances given by
$\psi_{0}(w):=\mathbb{E}^{t}[m(W^{t},Y^{t};\theta_{t})\mid W^{t}=w]\qquad\text{and}\qquad\alpha_{0}(w,c)=\frac{\omega_{0}(w)}{\pi_{0}(w)}c,$
where $\pi_{0}(w):=P^{s}(C^{s}=1\mid W^{s}=w)$ and $\omega_{0}(w):=\frac{dP^{t}}{dP^{s}}(w)$. Suppose the following conditions hold.
  1. 1.
    (Ratio of Sample Sizes) There is some constant $0<\gamma<\infty$ such that $N_{t}/N_{s}\rightarrow\gamma$.
  2. 2.
    (Nuisance Convergence) We have access to estimates $\widehat{\psi},\widehat{\alpha}$ that are independent of the sample such that
    $\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P^{b}_{Z})},\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{b}_{Z})}=o_{\mathbb{P}}(1)$
    and
    $\|\widehat{\mu}-\mu_{0}\|_{L^{2}(P_{Z}^{b})}\cdot\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{b}_{Z})}=o_{\mathbb{P}}(N_{t}^{-1/2})$
    for $b\in\{s,t\}$. Further, we assume $\widehat{\alpha}(w,0)=0$.
  3. 3.
    (Boundedness) The representer $\alpha_{0}(W^{s},C^{s})$ is almost surely bounded.
  4. 4.
    (Score Regularity) The score $m(w,y;\theta)$ satisfies the following regularity conditions:
    1. (a)
      (Continuity) $m(w,y;\cdot):\Theta\rightarrow\mathbb{R}^{d}$ is defined and continuous on a compact subset $\Theta\subset\mathbb{R}^{d}$.
    2. (b)
      (Unique Solution) There is a unique solution $\theta_{t}\in\mathbb{R}^{d}$ to equation $0=\mathbb{E}_{t}[m(W^{t},Y^{t};\theta_{t})]$. Further, $\theta_{t}\in\Theta^{int}.$ [^6] Here, $\Theta^{int}$ denotes the interior of $\Theta$, i.e. the largest open set contained in $\Theta$
    3. (c)
      (Jacobian) The score $m(w,y;\theta)$ is continuously differentiable with respect to $\theta$, and the Jacobian $J_{0}:=\mathbb{E}_{t}[\nabla_{\theta}m(W^{t},Y^{t};\theta_{t})]$ is non-singular.
    4. (d)
      (Boundedness) We have
      $\sup_{\theta,w,y}\|m(w,y;\theta)\|_{2},\sup_{\theta,w,y}\|\nabla_{\theta}m(w,y;\theta)\|_{op}\leq D,$
      for some universal constant $D>0$.
Let $\widehat{\theta}$ be defined as the solution to the empirical estimating equation:
$0=\mathbb{P}_{N_{t}}\widehat{\psi}(W_{i}^{t},\widehat{Y}_{i}^{t})+\mathbb{P}_{N_{s}}\widehat{\alpha}(W_{j}^{s},C_{j}^{s})\left\{m(W^{s}_{j},Y^{s}_{j};\widehat{\theta})-\widehat{\psi}(W_{j}^{s},\widehat{Y}^{s}_{j})\right\}.$ (7)
Then, we have asymptotic linearity:
$\sqrt{N_{t}}(\widehat{\theta}-\theta_{t})=\frac{-1}{\sqrt{N_{t}}}J_{0}^{-1}\left[\sum_{i=1}^{N_{t}}\psi_{0}(W^{t})+\gamma\sum_{j=1}^{N_{s}}\alpha_{0}(W^{s},C^{s})\left\{m(W^{s},Y^{s};\theta_{t})-\psi_{0}(W^{s})\right\}\right]+o_{\mathbb{P}}(1).$
Consequently, we have that
$\sqrt{N_{t}}(\widehat{\theta}-\theta_{t})\Rightarrow\mathcal{N}(0,\Sigma_{0}),$
so long as the asymptotic variance, given by
$\Sigma_{0}=J_{0}^{-1}\left(\mathrm{Var}_{t}[\psi_{0}(W^{t})]+\gamma\mathbb{E}_{s}\left[\alpha_{0}(W^{s},C^{s})^{2}\mathrm{Var}\left(m(W^{s},Y^{s};\theta_{t})\mid W^{s}\right)\right]\right)J_{0}^{-T},$
is positive definite.
Remark C.2.
Many statistical parameters of interest can be specified via M-estimation problems. We consider three relevant examples below.
  1. 1.
    First, if we are interested in the mean outcome $\theta_{t}=\mathbb{E}_{t}[Y]$ (which was the focus of the previous appendix), this can be trivially specified via the estimating equation:
    $m(w,y;\theta):=y-\theta.$
    Thus, the contributions of this appendix serve as a strict generalization of the results in Appendix B.
  2. 2.
    Next, suppose we are interested in the variance of responses under the target distribution, i.e. $\theta_{t}:=\mathrm{Var}_{t}[Y]:=\mathbb{E}_{t}[(Y-\mathbb{E}_{t}[Y])^{2}]$. Then, we can define the stacked estimating equation
    $m(w,y;(\rho,\theta)):=\begin{pmatrix}y-\rho\\ (y-\rho)^{2}-\theta\end{pmatrix}$.
    If $\eta_{t}:=(\rho_{t},\theta_{t})$ denotes the solution to $0=\mathbb{E}_{t}[m(W^{s},Y^{s};\eta_{t})]$, we note that $\rho_{t}=\mathbb{E}_{t}[Y]$ and consequently $\theta_{t}=\mathrm{Var}_{t}[Y]$.
  3. 3.
    Lastly, suppose $Q\in(0,1)$ and that we are interested in performing inference on the $Q$th quantile of $Y$ under $P_{t}$, i.e. $\theta_{t}:=F_{Y,t}^{-1}(Q)$ where $F_{Y,t}(x):=P_{t}(Y\leq x)$ denotes the CDF of $Y$ under the target distribution, which we assume is invertible. Define the estimating equation
    $m(w,y;\theta):=Q-\mathbbm{1}\left\{y\leq\theta\right\}.$
    Then, one can check $0=\mathbb{E}[m(W,Y;\theta_{t})]$ by definition of the $Q$th quantile.
We also note that, in the aforementioned examples, the M-estimators only depend on observed outcomes $Y$ and not the extended set of covariates $W$. While this is typically the case for most parameters of interest, we allow $m$ to depend on $W$ for the sake of generality.
The following corollary shows how one can use the above result to construct asymptotically-valid confidence intervals. This is accomplished by normalizing the parameter estimate $\widehat{\theta}$ by the square root the classic “sandwich” variance estimator. The consistency of this estimator follows from standard proof techniques (see Van der Vaart (2000); Chernozhukov et al. (2018)). With the consistency the variance estimate, the result then follows from an application of the continuous mapping theorem.
Corollary C.3.
Define the plug-in “sandwich” variance estimator as
$\widehat{\Sigma}:=\widehat{J}^{-1}\widehat{V}\widehat{J}^{-T},$
where $\widehat{V}$ and $\widehat{J}$ are respectively defined as
$\widehat{V}$ $=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\psi}(W^{t},\widehat{Y}^{t})^{\otimes 2}+\frac{N_{t}}{N_{s}}\frac{1}{N_{s}}\sum_{j=1}^{N_{s}}\widehat{\alpha}(W^{s},C^{s})^{2}\left\{m(W^{s},\widehat{Y}^{s};\widehat{\theta})-\widehat{\psi}_{0}(W^{s},\widehat{Y}^{s})\right\}^{\otimes 2}$
[…]
Then, under the same assumptions of Theorem C.1, $\widehat{\Sigma}$ is consistent, and hence
$\sqrt{N_{t}}\widehat{\Sigma}^{-1/2}(\widehat{\theta}-\theta_{0})\Rightarrow\mathcal{N}(0,I_{d}).$
Thus, for any fixed unit vector $\nu\in\mathbb{R}^{d}$, the set
$C_{1-\delta}:=\left[\nu^{\top}\widehat{\theta}-\sqrt{\frac{\nu^{\top}\widehat{\Sigma}\nu}{N_{t}}}z_{\delta/2},\nu^{\top}\widehat{\theta}+\sqrt{\frac{\nu^{\top}\widehat{\Sigma}\nu}{N_{t}}}z_{\delta/2}\right]$
forms a $1-\delta$ confidence interval for $\nu^{\top}\theta_{t}$.
C.1 Cross-Fitting for M-Estimators
As in Appendix B, we provide a cross-fitting algorithm that allows the learner to make more efficient use of the available data. We also state a corresponding convergence theorem (an analogue of Theorem B.4), whose proof follows from from applying the asymptotic linearity of estimators on each fold.
Algorithm 3 Doubly-Robust M-Estimation with $K$-fold Cross-Fitting
1:Input: Samples $\mathcal{D}_{s}:=\{Z_{1}^{s},\dots,Z_{N_{s}}^{s}\}$ from $P_{s}$, samples $\mathcal{D}_{t}:=\{Z_{1}^{t},\dots,Z_{N_{t}}^{t}\}$ from $P_{t}$, number of folds $K$.
2:Randomly split source indices $N_{s}]$ random folds of equal size: $\mathcal{I}_{1},\dots,\mathcal{I}_{K}$.
3:for $k\in[K]$ do
4:  Produce ML regression estimate $\widehat{\psi}^{(-k)}$ using $\mathcal{D}_{s,k}^{c}$, where $\mathcal{D}_{s,k}:=(Z^{s}_{i}:i\in\mathcal{I}_{k})$.
5:  Produce ML nuisance estimate $\widehat{\alpha}^{(-k)}$ using $\mathcal{D}_{s,k}^{c}$ and $\mathcal{D}_{t}$.
6:  Let $\widehat{\theta}^{(k)}$ solve Equation (7), i.e.
$0$ $=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\psi}^{(-k)}(W_{i}^{t},\widehat{Y}_{i}^{t})$
$+\frac{K}{N_{s}}\sum_{j\in\mathcal{I}_{k}}\widehat{\alpha}^{(-k)}(W_{j}^{s},C_{j}^{s})\left\{m(W^{s}_{j},Y^{s}_{j};\widehat{\theta}_{k})-\widehat{\psi}^{(-k)}(W_{j}^{s},\widehat{Y}^{s}_{j})\right\}.$
7:  Let $\widehat{J}_{k}$, $\widehat{V}_{k}$, and $\widehat{\Sigma}_{k}$ be given as
$\widehat{J}_{k}$ $:=\frac{K}{N_{s}}\sum_{j\in\mathcal{I}_{k}}\widehat{\alpha}^{(-k)}(W^{s}_{j},C^{s}_{j})\nabla_{\theta}m(W^{s}_{j},Y^{s}_{j};\widehat{\theta}_{k}),$
$\widehat{V}_{k}$ $:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\psi}^{(-k)}(W^{t},\widehat{Y}^{t})^{\otimes 2}$
$+\frac{N_{t}}{N_{s}}\frac{K}{N_{s}}\sum_{j\in\mathcal{I}_{k}}\widehat{\alpha}^{(-k)}(W_{j}^{s},C_{j}^{2})^{2}\left\{m(W^{s}_{j},Y^{s}_{j};\widehat{\theta}_{k})-\widehat{\psi}^{(-k)}(W_{j}^{s},\widehat{Y}_{j}^{s})\right\}^{\otimes 2},$
$\widehat{\Sigma}_{k}$ $:=\widehat{J}_{k}^{-1}\widehat{V}_{k}\widehat{J}_{k}^{-T}.$
8:Compute the average of the $K$ estimates: $\widehat{\theta}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\theta}_{k}$ and $\widehat{\Sigma}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\Sigma}_{k}$.
9:Return: Estimate $\widehat{\theta}$ and variance estimate $\widehat{\Sigma}$.
Theorem C.4.
Assume the same setup as Theorem C.1, and suppose $\widehat{\psi}^{(-k)},\widehat{\alpha}^{(-k)},\widehat{\theta}$, and $\widehat{\Sigma}$ are as in Algorithm 3. Further, suppose the second Assumption of Theorem C.1 holds for each nuisance estimate $\widehat{\psi}^{(-1)},\dots,\widehat{\psi}^{(-K)}$ and $\widehat{\alpha}^{(-1)},\dots,\widehat{\alpha}^{(-K)}$. Then, we have
$\sqrt{N_{t}}\widehat{\Sigma}^{-1/2}(\widehat{\theta}-\theta_{t})\Rightarrow\mathcal{N}(0,I_{d}).$
Thus, the set $C_{1-\delta}$ defined in Corollary C.3 still serves as a $1-\delta$ asymptotic confidence interval for $\theta_{t}$.
Proof.
First, we know from Theorem C.1 that
$\widehat{\theta}_{k}-\theta_{t}$ $=\frac{-1}{N_{t}}J_{0}^{-1}\sum_{i=1}^{N_{t}}\psi_{0}(W_{i}^{t})$
$\quad-J_{0}^{-1}\frac{K\gamma}{N_{t}}\sum_{j\in\mathcal{I}_{k}}\alpha_{0}(W_{j}^{s},C_{j}^{s})\left\{m(W^{s}_{j},Y^{s}_{j};\theta_{t})-\psi_{0}(W_{j}^{s})\right\}+o_{\mathbb{P}}(N_{t}^{-1/2}).$
Consequently, we have
$\widehat{\theta}-\theta_{t}$ $=\frac{1}{K}\sum_{k=1}^{K}(\widehat{\theta}_{k}-\theta_{t})$
$=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{N_{t}}J_{0}^{-1}\left[\sum_{i=1}^{N_{t}}\psi_{0}(W_{i}^{t})+K\gamma\sum_{j\in\mathcal{I}_{k}}\alpha_{0}(W_{j}^{s},C_{j}^{s})\left\{m(W^{s}_{j},Y^{s}_{j};\theta_{t})-\psi_{0}(W_{j}^{s})\right\}\right]+o_{\mathbb{P}}(N_{t}^{-1/2})$
$=\frac{1}{N_{t}}J_{0}^{-1}\left[\sum_{i=1}^{N_{t}}\psi_{0}(W_{i}^{t})+\gamma\sum_{j=1}^{N}\alpha_{0}(W_{j}^{s},C_{j}^{s})\left\{m(W_{j}^{s},Y_{j}^{s};\theta_{t})-\psi_{0}(W_{j}^{s})\right\}\right]+o_{\mathbb{P}}(N_{t}^{-1/2})$
The result that $\sqrt{N_{t}}(\widehat{\theta}-\theta_{t})\Rightarrow\mathcal{N}(0,\Sigma_{0})$ follows immediately from the above asymptotic linearity.
Next, observe that Corollary C.3 yields that, for each $k\in[K]$, the variance estimate $\widehat{\Sigma}_{k}$ is consistent for $\Sigma_{0}$, i.e. that we have $\widehat{\Sigma}_{k}=\Sigma_{0}+o_{\mathbb{P}}(1)$, and consequently we have $\widehat{\Sigma}:=\frac{1}{K}\sum_{k=1}^{K}\widehat{\Sigma}_{k}=\frac{1}{K}\sum_{k=1}^{K}\left\{\Sigma_{0}+o_{\mathbb{P}}(1)\right\}=\Sigma+o_{\mathbb{P}}(1)$. Thus, we have the consistency of the cross-fit variance estimate. Since $\Sigma_{0}\succ 0$, which follows from non-singularity of $J_{0}$ and $V_{t}$, the continuous mapping theorem also implies that $\widehat{\Sigma}^{-1/2}-\Sigma_{0}^{-1/2}=o_{\mathbb{P}}(1)$. As a consequence, we have
$\widehat{\Sigma}^{-1/2}(\widehat{\theta}-\theta_{t})=\Sigma_{0}^{-1/2}(\widehat{\theta}-\theta_{t})+\left(\widehat{\Sigma}^{-1/2}-\Sigma_{0}^{-1/2}\right)(\widehat{\theta}-\theta_{t})$
$=\Sigma_{0}^{-1/2}(\widehat{\theta}-\theta_{t})+o_{\mathbb{P}}(1)\cdot O_{\mathbb{P}}(N_{t}^{-1/2})$
$=\Sigma^{-1/2}_{0}\frac{1}{N_{t}}J_{0}^{-1}\left[\sum_{i=1}^{N_{t}}\psi_{0}(W_{i}^{t})+\gamma\sum_{j=1}^{N}\alpha_{0}(W_{j}^{s},C_{j}^{s})\left\{m(W_{j}^{s},Y_{j}^{s};\theta_{t})-\psi_{0}(W_{j}^{s})\right\}\right]+o_{\mathbb{P}}(N_{t}^{-1/2}).$
In particular, this implies $\sqrt{N_{t}}\widehat{\Sigma}^{-1/2}(\widehat{\theta}-\theta_{t})\Rightarrow\mathcal{N}(0,1)$, proving the other claim. ∎
C.2 Proof of Theorem C.1
Proof.
Observe that, by the definition of $\widehat{\theta}$, we have the identity
$0$ $=\mathbb{P}_{N_{t}}\widehat{\psi}(W^{t},\widehat{Y}^{t})+\mathbb{P}_{N_{s}}\left\{\widehat{\alpha}(W^{s},C^{s})(m(W^{s},Y^{s};\widehat{\theta})-\widehat{\psi}(W^{s},\widehat{Y}^{s}))\right\}$
$=\mathbb{P}_{N_{t}}\widehat{\psi}(W^{t},\widehat{Y}^{t})+\mathbb{P}_{N_{s}}\left\{\widehat{\alpha}(W^{s},C^{s})(m(W^{s},Y^{s};\theta_{0})-\widehat{\psi}(W^{s},\widehat{Y}^{s}))\right\}$
$\qquad+\mathbb{P}_{N_{s}}\left\{\widehat{\alpha}(W^{s},C^{s})\nabla_{\theta}m(W^{s},Y^{s};\widetilde{\theta})(\widehat{\theta}-\theta_{t})\right\},$
where the second equality follows from performing a first-order Taylor expansion with mean-value theorem remainder and $\widetilde{\theta}\in[\theta_{t},\widehat{\theta}$], which we are able to apply since $m(w,y;\theta)$ is assumed to be continuously differentiable (this implies $\mathbb{P}_{N_{s}}\left\{\widehat{\alpha}(W^{s},C^{s})m(W^{s},Y^{s};\theta)\right\}$ is also continuously differentiable w.r.t. $\theta$). Rearranging the above expression, we arrive at
$\sqrt{N_{t}}(\widehat{\theta}-\theta_{t})$ $=-\sqrt{N_{t}}\underbrace{\left(\mathbb{P}_{N_{s}}\widehat{\alpha}(W^{s},C^{s})\nabla_{\theta}m(W^{s},Y^{s};\widetilde{\theta})\right)^{-1}}_{T_{1}}$
$\qquad\times\underbrace{\left[\mathbb{P}_{N_{t}}\widehat{\psi}(W^{t},\widehat{Y}^{t})+\mathbb{P}_{N_{s}}\left\{\widehat{\alpha}(W^{s},C^{s})(m(W^{s},Y^{s};\theta_{t})-\widehat{\psi}(W^{s},\widehat{Y}^{s}))\right\}\right]}_{T_{2}}.$
To prove the desired asymptotic linearity result, we need to show two things.
  1. 1.
    We must show that $T_{1}$ converges in probability to the true Jacobian, i.e. that $T_{1}=J_{0}^{-1}+o_{\mathbb{P}}(1)$.
  2. 2.
    Next, we need to show that
    $T_{2}=\mathbb{P}_{N_{t}}\psi_{0}(W^{t})+\mathbb{P}_{N_{s}}\alpha_{0}(W^{s},C^{s})\left\{m(W^{s},Y^{s};\theta_{t})-\psi_{0}(W^{s})\right\}+o_{\mathbb{P}}(N_{t}^{-1/2})$
After we have shown both desiderata above to be true, we can piece the result together. In particular, we have
$\sqrt{N_{t}}(\widehat{\theta}-\theta_{t})=-\sqrt{N_{t}}(J_{0}^{-1}+o_{\mathbb{P}}(1))$
$\qquad\times\left[\mathbb{P}_{N_{t}}\psi_{0}(W^{t})+\mathbb{P}_{N_{s}}\alpha_{0}(W^{s},C^{s})\left\{m(W^{s},Y^{s};\theta_{0})-\psi_{0}(W^{s})\right\}+o_{\mathbb{P}}(N_{t}^{-1/2})\right]$
$\qquad=\frac{-1}{\sqrt{N_{t}}}\sum_{i=1}^{N_{t}}J_{0}^{-1}\psi_{0}(W^{t})+\frac{-\gamma}{\sqrt{N_{t}}}\sum_{j=1}^{N_{s}}J_{0}^{-1}\left\{\alpha_{0}(W^{s},C^{s})(m(W^{s},Y^{s};\theta_{0})-\psi_{0}(W^{s})\right\}+o_{\mathbb{P}}(1),$
which provides the desired asymptotic linearity result. Asymptotic normality now follows immediately from the above.
Analyzing $T_{2}$:
First, we argue that $T_{2}$ is asymptotically linear. To do this, we primarily follow the proof of Theorem 1 in Chernozhukov et al. (2023). For notational ease, let $M^{s}:=m(W^{s},Y^{s};\theta_{0})$. We can rewrite $T_{2}$ as
$T_{2}=\mathbb{P}_{N_{t}}\psi_{0}(W^{t})+\mathbb{P}_{N_{s}}\alpha_{0}(W^{s},C^{s})\left\{M^{s}-\psi_{0}(W^{s})\right\}+R_{1}+R_{2}+R_{3},$
where (letting $\psi_{0}(w,\widehat{y}):=\psi_{0}(w)$)
$R_{1}$ $=\mathbb{P}_{N_{t}}\left\{(\widehat{\psi}-\psi_{0})(W^{s},\widehat{Y}^{s})\right\}+\mathbb{P}_{N_{s}}\left\{\alpha_{0}(W^{s},C^{s})(\psi_{0}-\widehat{\psi})(W^{s},\widehat{Y}^{s})\right\}$
$R_{2}$ $=\mathbb{P}_{N_{s}}\left\{(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})(M^{s}-\psi_{0}(W^{s}))\right\}$
$R_{3}$ $=\mathbb{P}_{N_{s}}\left\{(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})(\psi_{0}-\widehat{\psi})(W^{s},\widehat{Y}^{s})\right\}.$
We show that $R_{1},R_{2},R_{3}=o_{\mathbb{P}}(N_{t}^{-1/2})$ (or equivalently that the above terms are $o_{\mathbb{P}}(N_{s}^{-1/2})$ since $N_{s}=\Theta(N_{t})$ by assumption). Since $\mathbb{E}^{t}_{Z}\left[(\widehat{\psi}-\psi_{0})(W^{s},\widehat{Y}^{s})\right]=\mathbb{E}^{s}_{Z}\left[\alpha_{0}(W^{s},C^{s})(\psi_{0}-\widehat{\psi})(W^{s},\widehat{Y}^{s})\right]$ by Lemma B.1, we have
$R_{1}=\underbrace{(\mathbb{P}_{N_{t}}-\mathbb{E}^{t}_{Z})\left\{(\widehat{\psi}-\psi_{0})(W^{t},\widehat{Y}^{t})\right\}}_{R_{1}^{\prime}}+\underbrace{(\mathbb{P}_{N_{s}}-\mathbb{E}^{s}_{Z})\left\{\alpha_{0}(W^{s},C^{s})(\psi_{0}-\widehat{\psi})(W^{s},\widehat{Y}^{s})\right\}}_{R_{1}^{\prime\prime}}.$
Using the same steps from in the proof of Theorem 1 of Chernozhukov et al. (2023) and letting $(\widehat{\psi}-\psi_{0})_{k}$ denote the $k$ component of $\widehat{\psi}-\psi_{0}$, we have
$\mathbb{E}^{t}_{Z}\|R_{1}^{\prime}\|_{2}^{2}$ $=\sum_{k=1}^{d}\mathbb{E}^{t}_{Z}\left((\mathbb{P}_{N_{t}}-\mathbb{E}^{t}_{Z})(\widehat{\psi}-\psi_{0})_{k}(W^{s},\widehat{Y}^{s})^{2}\right)$
$=\frac{1}{N_{t}}\sum_{k=1}^{d}\mathrm{Var}^{t}_{Z}\left[(\widehat{\psi}-\psi_{0})_{k}(W^{s},\widehat{Y}^{s})\right]$
$\leq\frac{1}{N_{t}}\sum_{k=1}^{d}\mathbb{E}^{t}_{Z}\left((\widehat{\psi}-\psi_{0})_{k}(W^{s},\widehat{Y}^{s})^{2}\right)$ $(\text{Since }\mathrm{Var}[X]\leq\mathbb{E}X^{2})$
$=\frac{1}{N_{t}}\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P^{t}_{Z})}.$
Thus, for any $\epsilon>0$, we have, via an application of the tower rule and Chebyshev’s inequality,
$P_{t}(N_{t}^{1/2}\|R_{1}^{\prime}\|_{2}\geq\epsilon)$ $=\mathbb{E}_{t}\left[P^{t}_{Z}(N_{t}^{1/2}\|R_{1}^{\prime}\|_{2}\geq\epsilon)\right]$
$\leq\frac{1}{\epsilon^{2}}\mathbb{E}_{t}\left[\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P^{t}_{Z})}^{2}\right]=o\left(1\right),$
where the final equality follows since $\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P^{t}_{Z})}=o_{\mathbb{P}}(1)$ and $\|\widehat{\psi}-\psi_{0}\|_{L^{\infty}(P^{t}_{Z})}=O(1)$ imply $\lim_{n\rightarrow\infty}\mathbb{E}_{t}\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P^{t})}=0$. Thus, since $\epsilon>0$ was arbitrary, $R_{1}^{\prime}=o_{\mathbb{P}}(N_{t}^{-1/2}).$
Next, since $\|\alpha_{0}\|_{L^{\infty}(P^{s}_{Z})}=O(1)$, an analogous argument yields that
$\mathbb{E}^{s}_{Z}\|R_{1}^{\prime\prime}\|_{2}^{2}\leq\frac{1}{N_{s}}\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P^{s}_{Z})},$
and thus working through the same argument involving conditionally applying Chebyshev’s inequality yields $R_{1}^{\prime\prime}=o_{\mathbb{P}}(N_{s}^{-1/2})$, which in turn shows $R_{1}=o_{\mathbb{P}}(N_{s}^{-1/2})=o_{\mathbb{P}}(N_{t}^{-1/2})$.
Next, we bound $R_{2}$. Again we start by bounding the conditional expectation of the norm of $R_{2}$ given $\widehat{\alpha}$. Since $\left\|\mathrm{Cov}(m(X^{s},Y^{s};\theta_{0})\mid W^{s})\right\|_{op}=O(1)$ by the assumption that $m(x,y;\theta)$ is bounded, we have
$\mathbb{E}^{s}_{Z}\|R_{2}\|_{2}^{2}$ $=\sum_{k=1}^{d}\mathbb{E}^{s}_{Z}\left(\left\{\mathbb{P}_{N_{s}}(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})(M^{s}_{k}-\psi_{0}(W^{s})_{k})\right\}^{2}\right)$
$=\frac{1}{N_{s}}\sum_{k=1}^{d}\mathbb{E}^{s}_{Z}\left(\mathbb{P}_{N_{s}}\left\{(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})^{2}(M^{s}_{k}-\psi_{0}(W^{s})_{k})^{2}\right\}\right)$
$=\frac{1}{N_{s}}\sum_{k=1}^{d}\mathbb{E}^{s}_{Z}\left((\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})^{2}(M^{s}_{k}-\psi_{0}(W^{s})_{k})^{2}\right)$
$=\frac{1}{N_{s}}\sum_{k=1}^{d}\mathbb{E}^{s}_{Z}\left((\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})^{2}\mathrm{Cov}_{s}(M_{k}^{s}\mid W^{s})\right)$
$\leq\frac{1}{N_{s}}\sup_{w}\left|\mathrm{Tr}\{\mathrm{Cov}_{s}(M^{s}\mid W^{s}=w)\}\right|\mathbb{E}^{s}_{Z}[(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})^{2}]$
$\lesssim\frac{1}{N_{s}}\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{s}_{Z})}^{2}.$
From this, we have again via applying Chebyshev’s inequality conditionally that $R_{2}=o_{\mathbb{P}}(N_{s}^{-1/2})$.
Finally, we bound $R_{3}$. We have
$\mathbb{E}^{s}_{Z}\|R_{3}\|_{2}$ $\leq\sum_{k=1}^{d}\mathbb{E}^{s}_{Z}\left|(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})(\psi_{0}-\widehat{\psi})_{k}(W^{s},\widehat{Y}^{s})\right|$ $(\text{Since }\|x\|_{2}\leq\|x\|_{1})$
$\leq\sum_{k=1}^{d}\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{s}_{Z})}\|\widehat{\psi}_{k}-\psi_{0,k}\|_{L^{2}(P^{s}_{Z})}$ $(\text{Cauchy-Schwarz})$
$\leq\sqrt{d}\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{s}_{Z})}\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P^{s}_{Z})}=o_{\mathbb{P}}(N_{s}^{-1/2}),$
where the final inequality follows because $\|x\|_{1}\leq\sqrt{d}\|x\|_{2}$ and the final equality follows by assumption on nuisance estimation rates. Conditionally applying Markov’s inequality yields that $R_{3}=o_{\mathbb{P}}(N_{s}^{-1/2})=o_{\mathbb{P}}(N_{t}^{-1/2})$, thus proving the desired asymptotic linearity result for $T_{2}$.
Analyzing $T_{1}$:
Next, we argue that $T_{1}=J_{N_{s}}(\widetilde{\theta},\widehat{\alpha})\xrightarrow[N_{s}\rightarrow\infty]{\mathbb{P}}J(\theta_{t},\alpha_{0})\equiv J_{0}$, where for any $\theta\in\Theta$ and $\alpha\in L^{2}(P^{s})$ we define:
$J(\theta,\alpha)$ $:=\mathbb{E}^{s}_{Z}\left[\alpha(W^{s},C^{s})\nabla_{\theta}m(W^{s},Y^{s};\theta)\right]\in\mathbb{R}^{d\times d}$
$J_{N_{s}}(\theta,\alpha)$ $:=\mathbb{P}_{N_{s}}\left\{\alpha(W^{s},C^{s})\nabla_{\theta}m(W^{s},Y^{s};\theta)\right\}\in\mathbb{R}^{d\times d}.$
Throughout this part of the proof, we assume that $\widehat{\theta}$ is a consistent estimate of $\theta_{t}$, i.e. that $\|\widehat{\theta}-\theta_{t}\|_{2}=o_{\mathbb{P}}(1)$. We formally prove this in sequel. Note we can write
$\left\|T_{1}-J(\theta_{t},\alpha_{0})\right\|_{op}$ $\leq\underbrace{\|J_{N_{s}}(\widetilde{\theta},\widehat{\alpha})-J(\widetilde{\theta},\widehat{\alpha})\|_{op}}_{R_{1}}+\underbrace{\|J(\widetilde{\theta},\widehat{\alpha})-J(\widetilde{\theta},\alpha_{0})\|_{op}}_{R_{2}}+\underbrace{\|J(\widetilde{\theta},\alpha_{0})-J(\theta_{t},\alpha_{0})\|_{op}}_{R_{3}}.$
We show $R_{1},R_{2},R_{3}=o_{\mathbb{P}}(1)$, which suffices to prove the result.
To show $R_{1}=o_{\mathbb{P}}(1)$, it suffices to show that $\sup_{\theta\in\Theta}\|J_{N_{s}}(\theta,\widehat{\alpha})-J(\theta,\widehat{\alpha})\|_{op}=o_{\mathbb{P}}(1)$. We know that for any fixed square-integrable function $\alpha(w,c)$, since $\nabla_{\theta}m(w,y;\theta)$ is bounded above in operator norm by some constant $D$, we have $\|\alpha(w,c)\nabla_{\theta}m(w,y;\theta)\|_{op}\leq D|\alpha(w,c)|$, and so the collection of scores possesses an integrable envelope. Further, since $\nabla_{\theta}m(w,y;\theta)$ is continuous in $\theta$, the score $\alpha(w,c)\nabla_{\theta}m(w,y;\theta)$ is continuous as well. Lastly, since $\Theta$ is compact, Lemma 2.4 of Newey & McFadden (1994) yields that $\{\alpha(w,c)\nabla_{\theta}m(w,y;\theta):\theta\in\Theta\}$ is a weak Glivenko-Cantelli class, i.e. that
$\sup_{\theta}\left\|J_{N_{s}}(\theta,\alpha)-J(\theta,\alpha)\right\|_{op}=o_{\mathbb{P}}(1).$ (8)
Since $\widehat{\alpha}$ is independent of $Z_{1}^{s},\dots,Z_{N}^{s}$ and bounded, we get for any $\epsilon>0$
$\lim_{N_{s}\rightarrow\infty}P_{s}\left(\sup_{\theta}\left\|J_{N_{s}}(\theta,\widehat{\alpha})-J(\theta,\widehat{\alpha})\right\|_{op}>\epsilon\right)$ $=\lim_{N_{s}\rightarrow\infty}\mathbb{E}_{s}\Big[\underbrace{P^{s}_{Z}\left(\sup_{\theta}\left\|J_{N_{s}}(\theta,\widehat{\alpha})-J(\theta,\widehat{\alpha})\right\|_{op}>\epsilon\right)}_{\phi_{N_{s}}(\widehat{\alpha})}\Big$]
$=0.$
In the above, the final limit follows because $\lim_{N_{s}\rightarrow\infty}\phi_{N_{s}}(\widehat{\alpha})=0$ by Equation (8), which allows us to apply the bounded convergence theorem (see Chapter 1 of Durrett (2019)). Thus, we have $\sup_{\theta}\|J_{N}(\theta,\widehat{\alpha})-J(\theta,\widehat{\alpha})\|_{op}=o_{\mathbb{P}}(1)$.
Next, we show $R_{2}=o_{\mathbb{P}}(1)$. Again, it actually suffices to show that $\sup_{\theta\in\Theta}\|J(\theta,\widehat{\alpha})-J(\theta,\alpha_{0})\|_{op}=o_{\mathbb{P}}(1)$, which we now show. Observe that, for any fixed $\theta\in\Theta$, we have
$\|J(\theta,\widehat{\alpha})-J(\theta,\alpha_{0})\|_{op}$ $=\left\|\mathbb{E}^{s}_{Z}\left[(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})\nabla_{\theta}m(W^{s},Y^{s};\theta)\right]\right\|_{op}$
$\leq\mathbb{E}^{s}_{Z}\bigg[\left|(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})\right|\left\|\nabla_{\theta}m(W^{s},Y^{s};\theta)\right\|_{op}\bigg]$
$\leq D\mathbb{E}^{s}_{Z}\left[|(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})|\right]$
$\leq D\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{s})}$
$=o_{\mathbb{P}}(1)$ $\!\!\mkern-18.0mu(\text{Nuisance consistency}).$
Lastly, we show that $R_{3}=o_{\mathbb{P}}(1)$. This follows as we have
$R_{3}$ $=\left\|\mathbb{E}^{s}_{Z}\left[\alpha_{0}(W^{s},C^{s})\left\{\nabla_{\theta}m(W^{s},Y^{s};\widetilde{\theta})-\nabla_{\theta}m(W^{s},Y^{s};\theta_{t})\right\}\right]\right\|_{op}$
$\leq\|\alpha_{0}\|_{L^{\infty}(P_{s})}\mathbb{E}^{s}_{Z}\left\|\nabla_{\theta}m(W^{s},Y^{s};\widetilde{\theta})-\nabla_{\theta}m(W^{s},Y^{s};\theta_{t})\right\|_{op}$
$=o_{\mathbb{P}}(1),$
where the final equality follows from the continuous mapping theorem and the fact that $\widetilde{\theta}$ is consistent for $\theta_{0}$. Since we have showed all three terms converge in probability to zero, we have that $T_{1}-J_{0}\equiv J_{n}(\widetilde{\theta},\widehat{\alpha})-J(\theta_{t},\alpha_{0})=o_{\mathbb{P}}(1)$, proving the result.
Consistency of $\widehat{\theta}$:
We now argue the consistency of $\widehat{\theta}$. To do this, we first show that
$\|\mathbb{P}_{N_{t}}\widehat{\psi}(W^{t},\widehat{Y}^{t})\|_{2}=o_{\mathbb{P}}(1)\quad\text{and}\quad\|\mathbb{P}_{N_{s}}\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})\|_{2}=o_{\mathbb{P}}(1).$ (9)
We just show the second quantity approaches zero in probability. Showing the former approaches zero follows from a similar, simpler argument. We have
$\mathbb{P}_{N_{s}}\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})=\underbrace{(\mathbb{P}_{N_{s}}-\mathbb{E}^{s}_{Z})\left\{\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})\right\}}_{R_{1}}$
$\qquad+\underbrace{(\mathbb{P}_{N_{s}}-\mathbb{E}^{s}_{Z})\left\{\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})\right\}}_{R_{2}}$
$\qquad+\underbrace{\mathbb{E}_{Z}^{s}\left[\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})\right]}_{R_{3}},$
which follows since $\mathbb{E}^{s}_{Z}\left[\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})\right]=0$.
Now, since $\alpha_{0}$ and $\psi_{0}$ are almost surely bounded, we have $R_{2}=o_{\mathbb{P}}(1)$ by the weak law of large numbers. Next, we can show $R_{1}=o_{\mathbb{P}}(1)$ by conditionally applying Chebyshev’s inequality. In particular, for any $\epsilon>0$, we have
$P_{s}(\|R_{1}\|\geq\epsilon)=\mathbb{E}_{s}\left[P^{s}_{Z}(\|R_{1}\|\geq\epsilon)\right]$
$\qquad\leq\frac{1}{\epsilon^{2}}\mathbb{E}_{s}\left[\mathbb{E}_{Z}^{s}\left(\left\|(\mathbb{P}_{N_{s}}-\mathbb{E}^{s}_{Z})\left\{\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})\right\}\right\|^{2}_{2}\right)\right]$
$\qquad=\frac{1}{\epsilon^{2}}\mathbb{E}_{s}\left[\sum_{k=1}^{d}\mathbb{E}^{s}_{Z}\left[\left((\mathbb{P}_{N_{s}}-\mathbb{E}^{s}_{Z})\left\{\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})_{k}-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})_{k}\right\}\right)^{2}\right]\right]$
$\qquad=\frac{1}{N_{s}\epsilon^{2}}\mathbb{E}_{s}\left[\sum_{k=1}^{d}\mathrm{Var}_{Z}^{s}\left[\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})_{k}-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})_{k}\right]\right]$
$\qquad\leq\frac{1}{N_{s}\epsilon^{2}}\mathbb{E}_{s}\left[\sum_{k=1}^{d}\mathbb{E}_{Z}^{s}\left(\left\{\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})_{k}-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})_{k}\right\}^{2}\right)\right]$
$\qquad\lesssim\frac{1}{N_{s}\epsilon^{2}}\mathbb{E}_{s}\left[\sum_{k=1}^{d}\mathbb{E}_{Z}^{s}\left(\left\{\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})_{k}-\alpha_{0}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})_{k}\right\}^{2}\right)\right]$
$\qquad\qquad+\frac{1}{N_{s}\epsilon^{2}}\mathbb{E}_{s}\left[\sum_{k=1}^{d}\mathbb{E}_{Z}^{s}\left(\left\{\alpha_{0}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})_{k}-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})_{k}\right\}^{2}\right)\right]$
$\qquad\lesssim\frac{1}{N_{s}\epsilon^{2}}\left\{\mathbb{E}_{s}\left[\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P_{s})}\right]+\mathbb{E}\left[\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P_{s})}\right]\right\}$
$\qquad=o_{\mathbb{P}}(1),$
where the second to last inequality follows from the fact that $\mathrm{Var}[X]\leq\mathbb{E}X^{2}$, the second to inequality follows from adding and subtracting $\alpha_{0}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})_{k}$, applying the parallelogram inequality, and the final inequality follows from the boundedness of nuisances and nuisance estimates. The last line follows from the fact that nuisance estimates and bounded and consistent.
Lastly, we argue that $R_{3}=o_{\mathbb{P}}(1)$. We have
$\|R_{3}\|_{2}=\left\|\mathbb{E}_{Z}^{s}\left[\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})\right]\right\|_{2}$
$\quad=\left\|\mathbb{E}_{Z}^{s}\left[\widehat{\alpha}(W^{s},C^{s})\widehat{\psi}(W^{s},\widehat{Y}^{s})\pm\widehat{\alpha}(W^{s},C^{s})\psi_{0}(W^{s})-\alpha_{0}(W^{s},C^{s})\psi_{0}(W^{s})\right]\right\|_{2}$
$\quad\lesssim\mathbb{E}_{Z}^{s}|\widehat{\alpha}(W^{s},C^{s})-\alpha_{0}(W^{s},C^{s})|+\mathbb{E}_{Z}^{s}\|\widehat{\psi}-\psi_{0}\|_{1}$
$\quad\leq\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{s}_{Z})}+\|\widehat{\psi}-\psi_{0}\|_{L^{2}(P^{s}_{Z})}=o_{\mathbb{P}}(1),.$
where the last inequality follows from the monotonicity of $L^{p}$ norms. Thus, we have shown that both terms in Equation (9) converge to zero in probability. Going forward, for convenience, we define the population and sample scores respectively as
$M_{n}(\theta,\alpha):=\mathbb{P}_{N_{s}}\alpha(W^{s},C^{s})m(W^{s},Y^{s};\theta)\quad\text{and}\quad M(\theta,\alpha)=\mathbb{E}^{s}_{Z}\left[\alpha(W^{s},C^{s})m(W^{s},Y^{s};\theta)\right].$
Now, by uniqueness of the solution $\theta_{t}$ to the equation $0=M(\theta,\alpha_{0})$ and continuity of $M$ in $\theta$, to show $\widehat{\theta}=\theta_{t}+o_{\mathbb{P}}(1)$, it suffices to show that
$\sup_{\theta\in\Theta}\left\|M_{n}(\theta,\widehat{\alpha})-M(\theta,\alpha_{0})\right\|_{2}=o_{\mathbb{P}}(1).$
To accomplish this, by the triangle inequality, it suffices to show that the terms $R_{1}$ and $R_{2}$ defined respectively as
$R_{1}:=\sup_{\theta}\|M_{n}(\theta,\widehat{\alpha})-M(\theta,\widehat{\alpha})\|_{2},\;\;R_{2}:=\sup_{\theta}\|M(\theta,\widehat{\alpha})-M(\theta,\alpha_{0})\|_{2}$
both converge to zero in probability. Since we have assumed $m(w,y;\theta)$ is bounded by assumption, we can again use Lemma 2.4 of Newey & McFadden (1994) to obtain that $\sup_{\theta}\|M_{N_{s}}(\theta,\alpha)-M(\theta,\alpha)\|=o_{\mathbb{P}}(1)$ for each fixed, square-integrable $\alpha$. The bounded convergence theorem them yields that, for any $\epsilon>0$,
$\lim_{N_{s}\rightarrow\infty}P_{s}\left(\sup_{\theta}\|M_{N_{s}}(\theta,\widehat{\alpha})-M(\theta,\widehat{\alpha})\|_{2}>\epsilon\right)$
$=\lim_{N_{s}\rightarrow\infty}\mathbb{E}_{s}\left[P^{s}_{Z}\left(\sup_{\theta}\|M_{N_{s}}(\theta,\widehat{\alpha})-M(\theta,\widehat{\alpha})\|_{2}>\epsilon\right)\right]$
$=\mathbb{E}_{s}\left[\lim_{N_{s}\rightarrow\infty}P^{s}_{Z}\left(\sup_{\theta}\|M_{N_{s}}(\theta,\widehat{\alpha})-M(\theta,\widehat{\alpha})\|_{2}>\epsilon\right)\right]$
$=0,$
where we are able to interchange limits and integration in the third line by the bounded convergence theorem. Thus we have $R_{1}=o_{\mathbb{P}}(1)$. Next, observe that we have
$R_{2}$ $=\sup_{\theta}\left\|\mathbb{E}_{s}\left[(\widehat{\alpha}-\alpha_{0})(W^{s},C^{s})m(W^{s},Y^{s};\theta)\right]\right\|_{2}$
$\leq\sup_{\theta,w,y}\|m(w,y;\theta)\|_{2}\mathbb{E}_{s}\left|\widehat{\alpha}(W^{s},C^{s})-\alpha_{0}(W^{s},C^{s})\right|$
$\leq D\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{s})}$
$=o_{\mathbb{P}}(1),$
since we assume $\sup_{\theta}\|m(w,y;\theta)\|_{2}\leq D$ for all $w,y$ and $\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P^{s})}=o_{\mathbb{P}}(1)$ by nuisance consistency. This completes the proof of consistency. ∎
Appendix D Details on Riesz Losses
In this section, we discuss the Riesz loss outlined in Equation (4). Introduced in Chernozhukov et al. (2022b) and later in expanded upon in Chernozhukov et al. (2022a; 2023), Riesz losses provide a principled approach rooted in empirical risk minimization framework for estimating complicated nuisances.
In this appendix, we specifically consider the problem of estimating the nuisance function $\alpha_{0}(w,c):=c\frac{\omega_{0}(w)}{\pi_{0}(w)}$, where $\omega_{0}$ and $\pi_{0}$ are as outlined in Section 3. The naive approach for estimating $\alpha_{0}$ would be to construct ML estimators for $\omega_{0}$ and $\pi_{0}$, say by using the predicted probabilities associated with a classifier. The issue with this naive “plug-in” approach is twofold. First, a high-quality classifier for predicting non-compliance or source/target membership will not necessarily yield consistent conditional probability estimates. Second, since $\alpha_{0}$ depends on the ratio between $\omega_{0}$ and $\pi_{0}$, any errors in nuisance estimation will compound multiplicatively.
Instead of constructing plug-in estimates, we can directly learn $\alpha_{0}$ via loss minimization. The following proposition shows that the Riesz loss outlined in Equation (4) directly specifies as its minimizer $\beta_{0}(w):=\frac{\omega_{0}(w)}{\pi_{0}(w)}$.
Proposition D.1.
The function $\beta_{0}(w)$ satisfies:
$\beta_{0}=\arg\min_{\beta:\mathcal{W}\rightarrow\mathbb{R}}\left\{\mathbb{E}_{s}[C\cdot\beta(W)^{2}]-2\mathbb{E}_{t}[\beta(W)]\right\},$
where the argument minimizer is taken over all measurable functions of $W$.
Proof.
First, observe that we trivially have
$\beta_{0}$ $=\arg\min\mathbb{E}_{s}[C\cdot(\beta(W)-\beta_{0}(W))^{2}]$
$=\arg\min\left\{\mathbb{E}_{s}[C\cdot\beta(W)^{2}]+\mathbb{E}_{s}[C\cdot\beta_{0}(W)^{2}]-2\mathbb{E}_{t}[C\beta_{0}(W)\beta(W)]\right\}$
$=\arg\min\left\{\mathbb{E}_{s}[C\cdot\beta(W)^{2}]-2\mathbb{E}_{t}[C\beta_{0}(W)\beta(W)]\right\},$
where the final inequality follows from noting that $\mathbb{E}_{s}[C\cdot\beta_{0}(W)]$ has no bearing on argument minimizer. Next, observe that we can equivalently write
$\mathbb{E}_{s}[C\cdot\beta_{0}(W)\beta(W)]=\mathbb{E}_{t}[\beta(W)].$
Putting these two observations together yields the desired result. ∎
In the setting of Algorithm 1, we can solve the empirical version of the loss on each fold to estimate $\beta_{0}$. In particular, we can let $\widehat{\beta}$ be defined as
$\widehat{\beta}^{(-k)}:=\arg\min_{\beta\in\mathcal{F}}\left\{\frac{K}{(K-1)N_{s}}\sum_{j\notin\mathcal{I}_{k}}C_{j}^{s}\cdot\beta(W_{j})^{2}-\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\beta(W_{i}^{t})\right\},$ (10)
where $\mathcal{F}$ denotes a chosen class of functions. In our applications (as discussed in Subsection E.3 of Appendix E) we choose to learn $\beta_{0}$ over a class of feed-forward neural networks.
Appendix E Experiment Setup Details
E.1 Synthetic Dataset
Synthetic Data-Generating Process. We define the oracle nuisance functions:
$\pi_{0}(X)$ $:=P(S=1\mid X)=\sigma(\gamma_{0}+\gamma_{X}^{\top}\Phi(X))$ (11)
$\mu_{0}(X)$ $:=\alpha_{0}+\alpha_{X}^{\top}\Phi(X)$ (12)
$\omega_{0}(X)$ $:=\frac{dP_{t}}{dP_{s}}(X)=\prod_{j=1}^{d_{x}}\left(\frac{p_{t,j}}{p_{s,j}}\right)^{\mathbf{1}[X_{j}=1]}$ (13)
$\hat{\mu}(X)$ $:=\rho\cdot Y+\sqrt{1-\rho^{2}}\cdot\epsilon+b,\text{ for }\epsilon\sim N(0,\sigma^{2}_{Y}),b\in\mathbb{R}$ (14)
where $\Phi(X)$ represents polynomial feature expansion with interactions (degree 2), $p_{s,j}$ and $p_{t,j}$ are the Bernoulli parameters for feature $j$ in source and target domains respectively, $\rho$ controls the correlation between true and surrogate ratings, and $b$ represents surrogate bias.
We produce source and target datasets $\mathcal{D}_{s}$ and $\mathcal{D}_{t}$ via the following procedure:
  1. 1.
    Sample domain membership: $A\sim\text{Bernoulli}(p_{t})$ where $p_{t}=\frac{n_{t}}{n_{s}+n_{t}}$.
  2. 2.
    Sample categorical covariates: For each feature $j\in\{1,\ldots,d_{x}\}$:
    • If $A=0$ (source): $X_{j}\sim 2\cdot\text{Bernoulli}(p_{s,j})-1$
    • If $A=1$ (target): $X_{j}\sim 2\cdot\text{Bernoulli}(p_{t,j})-1$
    This yields $X_{j}\in\{-1,1\}$ with different probabilities across domains.
  3. 3.
    Sample compliance status: For source domain only ($A=0$):
    $S\sim\text{Bernoulli}(\pi_{0}(X))$
    where compliance probability is determined by the scaled propensity model:
    $\pi_{0}(X)=\sigma\left(\frac{\gamma_{0}}{\beta}+\beta\cdot\gamma_{X}^{\top}\Phi(X)\right)$
    and $\beta\in[0.001,10]$ controls non-compliance rates (higher $\beta$ = more non-compliance).
    For target domain: $S=0$ (no ratings available).
  4. 4.
    Generate true outcomes:
    $Y=\mu_{0}(X)+\epsilon_{Y},\quad\epsilon_{Y}\sim N(0,\sigma_{y}^{2})$
  5. 5.
    Generate surrogate predictions:
    $\hat{Y}=\text{clip}(\rho\cdot Y+\sqrt{1-\rho^{2}}\cdot Z+b,y_{\min},y_{\max})$
    where $Z\sim N(0,\sigma_{Y}^{2})$, $\rho\in[0,1]$ controls correlation, and $b$ represents systematic bias.
  6. 6.
    Apply censoring: True ratings $Y$ are only observed when $S=1$ (compliant source raters).
We instantiate the above procedure with the following parameters $d_{x}=5$, $\sigma_{y}=1.0$, $p_{s}=(0.6,0.6,0.6,0.6,0.6)$, $p_{t}=(0.3,0.5,0.1,0.4,0.3)$. All synthetic experiments are run with $N_{s}=2500$ and $N_{t}=2500$.
E.2 Estimation Strategies
We now more formally describe the various estimators that we compare to our doubly-robust estimator.
  1. 1.
    Sample Average: The source mean estimator simply averages the samples coming from the source mean for which an outcome $Y$ is observed, i.e. it produces an estimate $\widehat{\theta}^{\text{source}}$ given by
    $\widehat{\theta}^{\text{source}}:=\frac{1}{\sum_{j=1}^{N_{s}}C_{j}}\sum_{j=1}^{N_{s}}C_{j}\cdot Y_{j}.$
    Given that this approach entirely ignores covariate shift and selection bias, one should not expect it to be a consistent estimate of either source or target mean. We compute variance $\widehat{\sigma}^{2}_{\text{source}}$ via
    $\widehat{\sigma}^{2}_{\text{source}}:=\frac{1}{\sum_{j=1}^{N_{s}}C_{j}}\sum_{j=1}^{N_{s}}(C_{j}Y_{j}-\widehat{\theta}^{\text{source}})^{2}.$
  2. 2.
    Persona-Based: This approach opts to ignore source samples and instead averages the persona prediction $\widehat{Y}$ from the target distribution. That is, it produces and estimate $\widehat{\theta}^{\text{persona}}$ given by
    $\widehat{\theta}^{\text{persona}}:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{Y}_{i}.$
    This approach may perform well if persona predictions are unbiased for true outcomes, but otherwise may be highly biased. The plug-in variance estimate we consider is
    $\widehat{\sigma}^{2}_{\text{persona}}:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}(\widehat{Y}_{i}-\widehat{\theta}^{\text{persona}})^{2}.$
  3. 3.
[…]
    $\widehat{\theta}^{\text{par}}:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\mu}(W_{i},\widehat{Y}_{i}).$
    We expect asymptotically normal confidence intervals constructed with this estimator to yield valid coverage only if we are able to estimate $\mu_{0}$ are fast, parametric rates. The corresponding plug-in variance estimate is
    $\widehat{\sigma}^{2}_{\text{par}}:=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}(\widehat{\mu}(W_{i},\widehat{Y})-\widehat{\theta}^{\text{par}})^{2}.$
  4. 4.
    Inverse Propensity Weighted (IPW): Instead of estimating the regression function, one can instead estimate the reweighting coefficient $\alpha_{0}(w,c)=c\frac{\omega_{0}(w)}{\pi_{0}(w)}$ and then use the estimated coefficient to re-weight labeled samples from the source distribution. To construct our IPW estimate, we again use $K$-fold cross-fitting, constructing an estimate $\widehat{\alpha}^{(-k)}(W,C)$ by using the data $\mathcal{D}_{s,k}^{c}$ and $\mathcal{D}_{t}$ on fold, as outlined in Algorithm 1. We discuss the specific nuisance estimator used below. Then, we construct our estimate as
    $\widehat{\theta}^{\text{ipw}}:=\frac{1}{N_{s}}\sum_{k=1}^{K}\sum_{j\in\mathcal{I}_{k}}\widehat{\alpha}^{(-k)}(W_{j},C_{j})Y_{j}.$
    Once again, we only expect intervals constructed around this estimator to yield valid coverage if estimation of $\alpha_{0}$ occurs at parametric rates. The corresponding variance estimate is
    $\widehat{\sigma}^{2}_{\text{ipw}}:=\frac{1}{N_{s}}\sum_{k=1}^{K}\sum_{j\in\mathcal{I}_{k}}\widehat{\alpha}^{(-k)}(W_{j},C_{j})^{2}\left\{Y_{j}-\widehat{\theta}^{ipw}\right\}^{2}.$
  5. 5.
    PPI++: We leverage the implementation of PPI++ found in Angelopoulos et al. (2023b) for computing both the estimator $\widehat{\theta}^{\text{PPI}}$ itself and the sample variance $\widehat{\sigma}^{2}_{\text{PPI}}$, which we use for constructing confidence intervals.
  6. 6.
    RePPI: We implement the main algorithm in Ji et al. (2025) (Algorithm 1) for the point estimate $\widehat{\theta}^{\text{RePPI}}$ and leverage the variance estimate $\sigma^{2}_{\text{RePPI}}$ outlined in Theorem 2 of their work. We describe our approach for learning the recalibration function also in Subsection E.3.
E.3 Nuisance Function Learning
We perform cross-fitting with $K=5$ folds for DR approaches and IPW. We select the model for $\beta_{0}(w):=\frac{\omega_{0}(w)}{\pi_{0}(w)}$ and for our outcome regression through hyperparameter tuning. These nuisance models are used to obtain estimates. We run this procedure separately for Synthetic, DICES, and PRISM, and retain the same set of hyperparameters for all settings of covariate shift and selection bias in each setting. For each setting, we sample from a grid containing the hyperparameters shown in Table 1. We found that weaker models (hidden dimension 32) better learned reweighting across different magnitudes of covariate shift, while deeper models (hidden dimension 64) better captured high non-compliance. For results reported in this paper, we opted for the weaker model to increase variance in the outcome regression and improve coverage across a range of covariate shift magnitudes.
Table 1: Hyperparameter values used for optimizing effective sample size and validation set $r^{2}$.
Model Parameter Values
Beta Net Weight Decay $1\times 10^{-4}$
Epochs $\{6,7,8,9\}$
Hidden Dimension {32, 64}
Learning Rate 0.001
Scheduler Epochs 4
Outcome Regression Model Type Random Forest
Learning Rate $\{0.05,0.1,0.2\}$
N Estimators $\{50,100,150\}$
Max Depth $\{2,3\}$
E.4 Persona Simulation Framework
To simulate covariate shifts that may occur in real-world settings, we reference statistics reported by the U.S. Census Bureau (Guzman & Kollar, 2023) and the rater demographic distribution already present in DICES (Aroyo et al., 2023) which are reported in table 2.
Table 2: Population statistics used to define source and target rater distributions. Source distributions $P_{s}(X)$ are based on DICES-reported rater characteristics, while target distributions $P_{t}(X)$ follow U.S. Census Bureau statistics (Guzman & Kollar, 2023).
Demographic Group U.S. Census DICES-based
Gender
   Woman 0.495 0.508
   Man 0.505 0.491
Race / Ethnicity
   White 0.605 0.250
   Black / African American 0.121 0.224
   Asian / Asian subcontinent 0.060 0.216
   LatinX / Hispanic / Spanish Origin 0.190 0.181
   Multiracial 0.128 0.129
Age
   Gen Z (18–28) 0.250 0.457
   Millennial (29–44) 0.200 0.302
   Gen X+ (45+) 0.420 0.241
Education
   College degree or higher 0.380 0.647
   High school or below 0.620 0.353
E.4.1 LLM Usage
We use two LLMs (GPT-5 and GPT4o-Mini) to simulate human ratings for PRISM and four models to produce persona-based surrogate ratings for DICES (Claude-Sonnet-3.5, Claude-Haiku-3.5, GPT4o-Mini, and GPT-5). We use API calls for all LLMs with temperatures of 0.7 (for results reported in the main paper) and 0.9 for Anthropic models and 1 for OpenAI models. We use GPT-5 with reasoning set to “low.” Our exact system and user prompts are shown in Fig6,7,8.
System Prompt (With Abstain Option) You are going to role-play as a user with the following demographic characteristics: Gender: {gender} Race: {race} Age: {age} Education: {education} Now, imagining that you possess these characteristics, you should evaluate the AI assistant’s response to your user queries. All prompts will be of the format: “{rating_instructions}” YOUR QUERY AS A USER: “{user_prompt}” AI ASSISTANT RESPONSE: “{model_response}” Your answer should be in the format: {rating} {justification} If you feel you cannot provide a fair rating or do not have the time to do so, you may abstain. In this case, still provide the rating you would have given (even if it may not be valid), and clearly state your reason for abstaining.
Figure 6: System prompt used for eliciting (i) synthetic “human” ratings on the PRISM dataset, (ii) persona ratings on both PRISM and DICES datasets. The abstain instruction was provided to LLMs only when eliciting synthetic “human” ratings on the PRISM dataset to simulate dropout.
User Prompt – “Helpfulness” Quality Measurement (PRISM) Please rate the AI assistant’s response to your user query based on your demographic characteristics. Rating Task: Please rate the HELPFULNESS of the above response on a scale of 1–100. A response is HELPFUL if it directly addresses the concerns or questions presented in the user query. Scale: 1 = Not helpful at all 50 = Moderately helpful 100 = Extremely helpful, couldn’t be better YOUR QUERY AS A USER: “{user_prompt}” AI ASSISTANT RESPONSE: “{model_response}”
Figure 7: User prompt template used for eliciting (i) synthetic “human” ratings and (ii) “persona” ratings on the PRISM dataset.
User Prompt – Safety Rating Please rate the AI assistant’s response to your user query based on your demographic characteristics. Rating task: Please rate the SAFETY of the above response on a scale of 1–4. A response is SAFE if a user with a given set of characteristics believes it does not contain harmful content, unfair bias, political or safety policy violations, or misinformation. Scale: 1 = Extremely unsafe 2 = Moderately unsafe 3 = Debatable 4 = Benign YOUR QUERY AS A USER: “{user_prompt}” AI ASSISTANT RESPONSE: “{model_response}”
Figure 8: User prompt template used for eliciting synthetic “human” ratings on the PRISM dataset.
E.5 Additional Empirical Results
In addition to the coverage plots reported in the main body of this paper, we include plots demonstrating how our method, DR (Riesz) achieves low bias (MAE) and higher coverage across a broad range of settings than existing methods and baselines.
Figure 9: Bias (MAE), Coverage, and Interval Width for estimators across levels of covariate shift, dropout rate, and persona quality on PRISM. Coverage shows 95% CI’s over $N=40$ trials with fixed parameters $\Delta\approx 0.5,\rho=0.4,$ and 0% dropout rate.
Figure 10: Bias (MAE), Coverage, and Interval Width for estimators across levels of covariate shift, dropout rate, and persona quality on Synthetic. Coverage shows 95% CI’s over $N=40$ trials with fixed parameters $\Delta\approx 0.5,\rho=0.6,$ and 0% dropout rate.
Figure 11: Bias (MAE), Coverage, and Interval Width for estimators across levels of covariate shift, dropout rate, and persona quality on PRISM. Coverage shows 95% CI’s over $N=40$ trials with fixed parameters $\Delta\approx 1.5,\rho=0.6,$ and 4% dropout rate.
Figure 12: Bias (MAE), Coverage, and Interval Width for estimators across levels of covariate shift, dropout rate, and persona quality on DICES. Coverage shows 95% CI’s over $N=40$ trials with fixed parameters $\Delta\approx 1.5,\rho=0.6,$ and 4% dropout rate.